### v1

# Hanke-Raus heuristic rule for iteratively regularized stochastic gradient descent

Abstract. In this work, we present a novel variant of the stochastic gradient descent method termed as iteratively regularized stochastic gradient descent (IRSGD) method to solve nonlinear ill-posed problems in Hilbert spaces. Under standard assumptions, we demonstrate that the mean square iteration error of the method converges to zero for exact data. In the presence of noisy data, we first propose a heuristic parameter choice rule (HPCR) based on the method suggested by Hanke and Raus, and then apply the IRSGD method in combination with HPCR. Precisely, HPCR selects the regularization parameter without requiring any a-priori knowledge of the noise level. We show that the method terminates in finitely many steps in case of noisy data and has regularizing features. Further, we discuss the convergence rates of the method using well-known source and other related conditions under HPCR as well as discrepancy principle. To the best of our knowledge, this is the first work that establishes both the regularization properties and convergence rates of a stochastic gradient method using a heuristic type rule in the setting of infinite-dimensional Hilbert spaces. Finally, we provide the numerical experiments to showcase the practical efficacy of the proposed method.
Keywords: Stochastic gradient descent; Iterative Landweber method; Convergence rates; Variational inequalities; Inverse problems; Nonlinear ill-posed problems; Regularization.
1. Introduction
This work focuses on deriving an approximate solution for a system of nonlinear ill-posed equations of the form
((1.1)) $F_{i}(u)=y^{\dagger}_{i},\hskip 5.69054pti=0,1,\ldots,P-1,$
[…]
((1.2)) $F(u)=y^{\dagger},$
where $F:U\rightarrow Y^{P}$ is defined as
$F(u)=\begin{pmatrix}F_{0}(u)\\ \vdots\\ F_{P-1}(u)\end{pmatrix}\ \ \text{and}\ \hskip 5.69054pty^{\dagger}=\begin{pmatrix}y^{\dagger}_{0}\\ \vdots\\ y^{\dagger}_{P-1}\end{pmatrix}.$
Usually, exact data $y^{\dagger}$ is never precisely available. Instead, a noisy data $y^{\delta}\in Y^{P}$ is available such that
((1.3)) $\|y^{\dagger}-y^{\delta}\|\leq\delta,$
for noise level $\delta\geq 0$. Here, we write $\delta=\sqrt{\delta_{0}^{2}+\delta_{1}^{2}+\cdots+\delta_{P-1}^{2}}$ and $y^{\delta}=(y^{\delta}_{0},y^{\delta}_{1},...,y^{\delta}_{P-1})$, where
((1.4)) $\|y_{i}^{\delta}-y_{i}^{\dagger}\|\leq\delta_{i}\ \ \text{for}\ \ i=0,1,2,\cdots,P-1.$
Due to ill-posedness of (1.1), its solution may not exist and even if it exists, it may not be unique. Moreover, the solution(s) can be highly sensitive to perturbations in the noisy data $y^{\delta}$ (see [6]). Consequently, regularization methods are needed for deducing the stable approximate solution of (1.1).
One of the prominent and highly effective classes of regularization methods is that of iterative regularization. This approach has achieved significant success in solving a wide variety of inverse problems (see [6, 16] and the references therein). Among these methods, the Landweber Iteration Method (LIM) is a well-studied and widely recognized classical iterative technique. For a Fréchet differentiable (or smooth) forward mapping $F:U\rightarrow Y$, LIM can be written as
((1.5)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta}),\hskip 28.45274ptk\geq 0.$
Here $u_{0}^{\delta}:=u_{0}$represents an initial guess that incorporates the prior knowledge about the solution to be reconstructed and $F^{\prime}(u_{k}^{\delta})^{*}$ is the adjoint of the Fréchet derivative of $F$ at $u_{k}^{\delta}$.
Scherzer [26] introduced a modification to the method (1.5) by introducing a damping term and named it as iteratively regularized Landweber iteration method. This method can be written as
((1.6)) $u_{k+1}^{\delta}=u_{k}^{\delta}-F^{\prime}(u_{k}^{\delta})^{*}(F(u_{k}^{\delta})-y^{\delta})-\lambda_{k}(u_{k}^{\delta}-u^{(0)}),\hskip 28.45274ptk\geq 0,$
where $\lambda_{k}\geq 0$. In comparison to the method (1.5), the convergence rates analysis of the method (1.6) needs fewer assumptions on the mapping $F$. Furthermore, the method (1.6) converges to a solution which is closed to $u^{(0)}$, however, the method (1.5) needs additional assumptions for the similar convergence behaviour (see [16] for nice details). This means that the inclusion of the additional damping term in (1.6) has several advantages over the method (1.5). When the noisy data is available, a proper stopping rule (a-posteriori) must be employed in order to show that the iterative schemes (1.5) and (1.6) are a regularization one. The commonly incorporated stopping rule is well known as discrepancy principle, i.e., the method is halted after $k_{\delta}=k_{\delta}(\delta,y^{\delta})$ steps, where
((1.7)) $\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|\leq\tau\delta<\|F(u_{k}^{\delta})-y^{\delta}\|,\hskip 28.45274pt0\leq k<k_{\delta},$
for some $\tau>1.$ In real-world applications, noise level is not always available or trustworthy, which makes the discrepancy principle (1.7) difficult to use. Consequently, while applying the discrepancy principle, overestimation or underestimation of the noise level may result in significant degradation of the reconstruction accuracy. Therefore, it is reasonable to work on heuristic stopping rules that do not rely on the knowledge of noise level. Real [25] proposed a heuristic rule for the Landweber iteration (1.5) in Banach space setting based on the discrepancy principle, by taking motivation from the method of Hanke and Raus [9]. For (1.5), heuristic rule determines an integer $k_{*}:=k_{*}(y^{\delta})$ by minimizing
((1.8)) $\Psi(k,y^{\delta}):=(k+a)\|F({u_{k}^{\delta}})-y^{\delta}\|^{p}$
over $k$, where $a\geq 1$ is a fixed number and $p>0$ is a constant.
It can be observed that for the conventional iterative regularization methods for solving inverse problems (such as the Landweber method [16], Levenberg-Marquardt method [5], conjugate gradient method [16], non-stationary Tikhonov iterative regularization method [20] and iteratively regularized Gauss-Newton method [16], among others), one of the common challenges is the significant computational burden at each iteration, mainly due to the need to process all the data, which can be excessively large and resource-intensive. To address this issue, a potential and effective approach is the utilization of stochastic gradient descent (SGD), a method introduced by Robbins et al. [24]. Further, Jin et al. [13] studied the convergence analysis of SGD method and demonstrated its regularizing nature when there is a noise in the data. The basic version of SGD can be written as (see [14])
((1.9)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\mu_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}),\hskip 28.45274ptk\geq 0,$
where $\mu_{k}$ is the corresponding step size and $i_{k}$ is the uniformly drawn index from the index set $\{0,1,...,P-1\}.$ It can be noted that the method (1.9) is a randomized version of (1.5). For (1.9), the stopping index $k(\delta)\in\mathbb{N}$ is defined using the following a-priori stopping rule
((1.10)) $\lim_{\delta\rightarrow 0^{+}}k(\delta)=\infty\hskip 14.22636pt\text{and}\hskip 14.22636pt\lim_{\delta\rightarrow 0^{+}}\delta^{2}\sum_{i=1}^{k(\delta)}\mu_{i}=0.$
We refer to [14, 17] for some recent literature on SGD method using a-priori stopping rule for linear as well nonlinear inverse problems and [12, 27] on SGD method using a-posteriori stopping rule for linear inverse problems. It is evident from these references that they exclusively use an a-priori or a-posteriori stopping rule in developing their convergence theories. From this observation, it is of significant interest to develop heuristic rules for stochastic gradient descent (SGD) type methods. Therefore, the primary objective of this paper is to explore the convergence analysis of the following iteratively regularized stochastic gradient descent (IRSGD) under heuristic stopping rules:
((1.11)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
$\hskip 85.35826pt\text{with}\quad 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2},$
where $\omega_{k}$ is the step size, $\lambda_{k}$ is the weighted parameter, $F^{\prime}_{i}(u)$ is the Fréchet derivative of $F_{i}(u)$, $i_{k}$ is an index randomly drawn from the index set $\{0,1,\ldots,P-1\}$ through uniform distributation. This method differs from the SGD method in the sense that an extra damping term $-\lambda_{k}(u_{k}^{\delta}-u_{0})$ is considered in its formulation. A key advantage of studying IRSGD lies in its ability to reflect SGD by setting the damping factor $\lambda_{k}=0$. Our approach is novel in the realm of stochastic methods for solving ill-posed inverse problems as it provides convergence analysis without relying on an a-priori stopping rule. Additionally, it circumvents the need to know the noise level, which further enhances its practical applicability.
For SGD method, Jin et al. [14] established the convergence rate in the mean squared norm in the context of nonlinear ill-posed problems, assuming an appropriate source condition. Further, Jahn et al. [12] further extended these results using an a-posteriori stopping rule, however, precise estimates for the convergence rates are not provided in that work. Inspired by these findings, one of the main aims of this work is to develop convergence rates for the iterates generated by (1.11) under an a-posteriori stopping rule. In addition to that, we also derive convergence rates under heuristic stopping criteria. To the best of our knowledge, this is the first study that addresses convergence rates of a stochastic type method for solving nonlinear ill-posed problems under heuristic stopping rule.
The main contributions of this paper are outlined as follows:
  • Instead of focusing on the analysis of a single equation, we turn our attention to exploring stochastic methods for solving systems comprising a large number of equations (i.e., $P$ is large in (1.1)). Stochastic methods involve the random consideration of each equation within equation (1.1) independently, thereby significantly reducing memory requirements. This approach finds relevance in numerous real-world scenarios (see [7, 21, 22]).
  • The method is fully data-driven, as it is applicable without using a-priori or a-posteriori stopping rule. This is the first attempt of construction of a stochastic method using heuristic type stopping rule. We establish the (strong) convergence of IRSGD method, which improves the weak convergence (i.e., convergence in probability) results provided in [12].
  • We establish the convergence rates for the proposed scheme (1.11) using a step size $\omega_{k}\in[\omega,\Omega]$ and a decaying weight parameter, under two types of stopping criteria: the discrepancy principle and a heuristic rule. The rates are comparable to those provided in [14, Theorem 2.5], [8, Theorem 3.2], and [13, Theorem 2.2] (see Table 1).
  • This work presents experimental evidence showcasing the method’s enhanced efficiency and accelerated performance achieved through the inclusion of a damping factor in SGD. This modification yields significantly better results compared to the standard SGD approach discussed in [14].
The rest of the paper is structured as follows. In Section 2, we list the preliminaries needed in our work. Section 3 covers the convergence analysis of our method in the presence of noisy and exact data. Section 4 discusses the convergence rates of our method for a-posteriori and heuristic stopping rules under certain assumptions. Section 5 presents numerical experiments on a system of ill-posed problems, demonstrating the applicability of the proposed scheme with heuristic stopping rule. The final section offers concluding remarks. Additionally, the Appendix includes few auxiliary lemmas required for our analysis.
2. Preliminaries
In this section, we review certain basic definitions and notation related to our work, see [2, 4, 14, 15, 23] for more details.
The notation $u_{k}$ is used to denote the iterates of our method corresponding to the exact data $y^{\dagger}.$ Additionally, we refer $\mathcal{F}_{k}$ as the filtration generated by the random indices $\{i_{1},...,i_{k-1}\}$ up to the $(k-1)$th iteration. Because we randomly select the indices $i_{k}$, the iteration $u_{k}^{\delta}$ in the IRSGD method becomes a random process. To assess its convergence, multiple methods are available. In our case, we will use the mean squared norm defined by $\mathbb{E}[\|\cdot\|^{2}]$, where $\mathbb{E}[\cdot]$ represents the expectation taken with respect to the filtration $\mathcal{F}_{k}$. It is important to note that the iterates $u_{k}^{\delta}$ are measurable with respect to $\mathcal{F}_{k}$ (see [15] for more details).
The following well-known identities will be frequently utilized: for any random variables $\Theta,\Theta_{1},\Theta_{2}$ and $\Phi$, and real constants $c_{1},c_{2}$, we have
  • (a)
    $\mathbb{E}[\Theta]=\mathbb{E}[\mathbb{E}[\Theta|\mathcal{F}_{k}]].$
  • (b)
    Additive: $\mathbb{E}[\Theta_{1}+\Theta_{2}|\mathcal{F}_{k}]=\mathbb{E}[\Theta_{1}|\mathcal{F}_{k}]+\mathbb{E}[\Theta_{2}|\mathcal{F}_{k}].$
    Linearity: $\mathbb{E}[c_{1}\Theta+c_{2}|\mathcal{F}_{k}]=c_{1}\mathbb{E}[\Theta|\mathcal{F}_{k}]+c_{2}.$
  • (c)
    If $\Theta\leq\Phi$, then $\mathbb{E}[\Theta|\mathcal{F}_{k}]\leq\mathbb{E}[\Phi|\mathcal{F}_{k}]$ almost surely.
Here the expression $\mathbb{E}[\Theta|\mathcal{F}_{k}]$ represents the expected value of $\Theta$ given the information contained in $\mathcal{F}_{k}$.
Triangular inequality for expectation: Let $\Theta_{1}$ and $\Theta_{2}$ be random variables. Then, there holds
((2.1)) $\mathbb{E}[\|\Theta_{1}+\Theta_{2}\|]\leq\mathbb{E}[\|\Theta_{1}\|]+\mathbb{E}[\|\Theta_{2}\|].$
Jensen’s inequality: For any random variable $\Theta$ and for any convex function $f$, there holds
((2.2)) $f(\mathbb{E}[\|\Theta\|])\leq\mathbb{E}[f(\|\Theta\|)],$
provided $\mathbb{E}[\|\Theta\|]<\infty$ and $\mathbb{E}[f(\|\Theta\|)]<\infty$ (see [1] for more details).
Next, we introduce an important result that splits the mean squared error $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]$, where $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ represents the error for any iteration $u_{k}^{\delta}$. This decomposition is a fundamental strategy in error analysis and plays a crucial role in deriving the convergence rates for iterative methods.
Bias-variance decomposition: The bias-variance decomposition for the iterative scheme is given by
((2.3)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}],$
where $B$ is defined as in (4.1) and $h\geq 0.$ In this decomposition:
  • The first term, $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}$, represents the bias, which includes both the approximation error and the error due to noise in the data.
  • The second term, $\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}]$, represents the variance, which accounts for the randomness introduced by the selection of the index $i_{k}$ during the iterations.
The bias reflects the deterministic error from the method’s approximation, while the variance captures the stochastic error due to the randomized nature of the iterative process.
We end this section by discussing few notations.
For some $u\in U$ and $\rho>0,$ we denote the open and closed balls in $U$ of radius $\rho$ centered at $u$, respectively by $B_{\rho}(u)$ and $\overline{B}_{\rho}(u).$ Let $\mathcal{D}_{\rho}(u^{\dagger})$ denote the set of all the solutions of (1.2) in $\overline{B}_{\rho}(u^{\dagger})$, i.e., $\mathcal{D}_{\rho}(u^{\dagger}):=\{u\in\overline{B}_{\rho}(u^{\dagger}):F(u)=y^{\dagger}\}$.
3. Iteratively regularized Stochastic gradient descent in Hilbert spaces
This section covers the convergence analysis of IRSGD method under certain assumptions. Precisely, this section is organised as follows. The subsection 3.1 aims to discuss the assumptions required for studying the convergence analysis. In subsection 3.2, we discuss a heuristic rule based on the modified discrepancy principle. Convergence of non-noisy iterates is discussed in subsection 3.3. Lastly, subsection 3.4 aims to demonstrate the convergence analysis of noisy iterates.
We recall that the IRSGD method can be expressed as
((3.1)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}),\hskip 14.22636ptk\geq 0,$
with $\text{with}\ 0\leq\lambda_{k}\leq\lambda_{\text{max}}<\frac{1}{2}$, where $F^{\prime}_{i_{k}}$ is the Frećhet derivative of $F_{i_{k}}$, $u_{0}^{\delta}$ is the initial guess of the exact solution, $\omega_{k}$ is the step size and the index $i_{k}$ is drawn randomly from the index set $\{0,1,\ldots,P-1\}$ at each iteration step.
3.1. Assumptions
In our work, we need the following assumptions.
Assumption 3.1.
For $i\in\{0,1,\ldots,P-1\}$ and $\rho>0$, there holds:
  • (i)
    The operator $F:U\rightarrow Y^{P}$ is continuous.
  • (ii)
    There exists a constant $L>0$ such that
    ((3.2)) $\|F_{i}^{\prime}(u)\|\leq L\ \text{for all}\ u\in B_{\rho+c(\rho)}(u_{0}),$
    where $L\leq\frac{1}{2\omega}$ (see Lemma 3.1 for the meaning of $c(\rho)$ and $\omega$). Moreover there is a constant $10<\eta<$ 1 such that
    ((3.3)) $\|F_{i}(\overline{u})-F_{i}(u)-F_{i}^{\prime}(u)(\overline{u}-u)\|\leq\eta\|F_{i}(\overline{u})-F_{i}(u)\|$
    for all $\overline{u},u\in B_{\rho+c(\rho)}(u_{0}).$
  • (iii)
    (1.2) has a solution $\hat{u}\in B_{\rho}(u_{0}).$
Motivated by [9], we consider the following heuristic rule for IRSGD in Hilbert spaces.
Rule 1.
Let $a\geq 1$ be a fixed number and let
$\Psi(k,y^{\delta}):=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}.$
We define $k_{*}:=k_{*}(y^{\delta})$ to be an integer such that
$k_{*}\in\arg\min\{\Psi(k,y^{\delta}):0\leq k\leq k_{\infty}\},$
where $k_{\infty}:=k_{\infty}(y^{\delta})$ is the largest integer such that $u^{\delta}_{k}\in\mathfrak{D}(F)$ for all $0\leq k<k_{\infty}$.
Remark 3.1.
Rule 1 is straightforward to implement in practice, as detailed in Section 5. During the iteration process, the value of $\Psi(k,y^{\delta})$ is monitored as a function of $k$. After a sufficient number of iterations, the process is terminated and $k_{*}$ is selected that minimizes $\Psi(k,y^{\delta})$. However, an important challenge associated with the nonlinear Landweber iteration [11] lies in determining an appropriate upper bound on the number of iterations. This is particularly critical due to its local convergence properties, which can lead to premature stopping or inaccurate identification of $k_{*}$. Therefore, the limit on the upper bound must be sufficiently large to avoid inaccurately estimating $k_{*}$ at the first local minimum, which is often not the global minimum [11]. To prevent $k_{*}$ from becoming too small, we have performed several numerical experiments on different values of $a$ in Rule 1.
Note that in order to prove Lemma 3.1 for IRSGD, it is essential to assume that $a\geq 1$. This Lemma will be further required to show the convergence of the scheme. Using the integer $k_{*}:=k_{*}(y^{\delta})$ determined by Rule 1, we call $u^{\delta}_{k_{*}(y^{\delta})}$ as an approximate solution.
A natural question arises concerning a family of noisy data $\{y^{\delta}\}$, where $y^{\delta}\to y^{\dagger}$ as $\delta\to 0$ is that how to guarantee that $u^{\delta}_{k_{*}(y^{\delta})}$ converges to a solution of (1.2) as $\delta\to 0\mathord{?}$
In general, the answer is no, as stated by Bakushinskii’s veto [3], which asserts that heuristic rules cannot guarantee the convergence in the worst-case scenario for any regularization method. Later, in [9], it is shown that the following assumption guarantee the convergence. We also need this condition in our work.
Assumption 3.2 (Noise condition).
The family of noisy data $\{y^{\delta}\}$ satisfies the condition $0<\|y^{\delta}-y^{\dagger}\|\rightarrow 0$ as $\delta\rightarrow 0$. Moreover, there exists a constant $\varkappa>0$ such that for every $y^{\delta}$ and for each $u\in S(y^{\delta})$, where $S(y^{\delta}):=\{u^{\delta}_{k}:0\leq k\leq k_{\infty}\}$ is constructed using (3.1), there holds
$\|y^{\delta}-F(u)\|\geq\varkappa\|y^{\delta}-y^{\dagger}\|.$
Under Assumption 3.2, it is immediate to see that Rule 1 comes out with a finite integer $k_{*}$. Indeed, if $k_{\infty}$ is finite, then there is nothing to prove. So we only consider that $k_{\infty}=\infty$. It then follows from Assumption 3.2 that
$\Psi(k,y^{\delta})=(k+a)\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\geq(k+a)\varkappa\|y^{\dagger}-y^{\delta}\|^{2}\rightarrow\infty$
as $k\rightarrow\infty$. Consequently, we can conclude that a finite integer $k_{*}$ exists, which achieves the minimum of $\Psi(k,y^{\delta})$.
3.2. Analysis for Rule 1
This subsection aims to demonstrate that
((3.4)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\rightarrow 0\quad\text{as}\quad\delta\rightarrow 0.$
Additionally, we demonstrate that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ under certain conditions.
In order to show (3.4), we propose an auxiliary integer $\hat{k}_{\delta}$ defined by the stopping rule
((3.5)) $\|F(u^{\delta}_{\hat{k}_{\delta}})-y^{\delta}\|+\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|<\|F(u^{\delta}_{k})-y^{\delta}\|+\frac{M}{k+a}\quad$
for $0\leq k<\hat{k}_{\delta}$, where $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)},\varrho>0$ is a constant, $\Omega>0,\eta$ is same as in (3.3), $c(\rho)$ is defined in (3.6) and $\tau>1$ is a real number (not to be confused with the one in (1.7)). Adding an extra term in the classical discrepancy principle produces a minor variation of the discrepancy principle (1.7). An essential component in establishing (3.4) is provided by this additional term, which ensures that $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0$.
Lemma 3.1.
Let Assumption 3.1 hold and let $\Omega,\omega$ be positive constants such that $\Omega\geq\omega$. Moreover, let $\kappa\in(0,1)$ and $\varrho>0$ be a sufficiently small constant and
((3.6)) $c(\rho):=\rho\lambda_{\textbf{max}}\frac{(1-\lambda_{\textbf{max}})+\sqrt{(1-\lambda_{\textbf{max}})^{2}+(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})(1+\kappa^{-2}\Omega^{2}L^{2})}}{(\frac{1}{1+\varrho}-(1-\lambda_{\text{max}})^{2})}.$
Then, under the following conditions
((3.7)) $D:=2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}-\tau^{-1}\Omega(2+M)(1+\eta)>0,$
((3.8)) $\sum_{k=0}^{\infty}\lambda_{k}\leq\lambda<\infty,$
there holds
  • (a)
    the stopping index $\hat{k}_{\delta}$ defined in (3.5) is finite for $\delta>0$ and $\hat{k}_{\delta}\rightarrow\infty$ as $\delta\rightarrow 0.$
  • (b)
    for a solution $\hat{u}$ to (1.2) and for any $0\leq k<\hat{k}_{\delta}$, if $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$, then $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0}).$
Proof.
First we prove assertion (b). For $k=0$, result trivially follows from (iii) of Assumption 3.1. By assuming that $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<\hat{k}_{\delta}$, we demonstrate that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u}).$ From (3.1), we have
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $=\|u_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0})-\hat{u}\|^{2}$
$=(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}\|u_{0}-\hat{u}\|^{2}+2\lambda_{k}(1-\lambda_{k})\langle u^{\delta}_{k}-\hat{u},u_{0}-\hat{u}\rangle$
((3.9)) $\hskip 14.22636pt+\omega_{k}^{2}\|F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\|^{2}+2\omega_{k}(1-\lambda_{k})\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F^{\prime}_{i_{k}}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$
$\hskip 14.22636pt+2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle.$
To estimate 3.2, we can write the last two terms of the sum as
$\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle$ $=\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),y_{i_{k}}^{\delta}-y_{i_{k}}^{\dagger}\rangle-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
((3.10)) $\hskip 5.69054pt-\langle y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}),F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(\hat{u})-F_{i_{k}}^{\prime}(u^{\delta}_{k})(u^{\delta}_{k}-\hat{u})\rangle,$
$2\omega_{k}\lambda_{k}\langle u_{0}-\hat{u},F^{\prime}_{i_{k}}(u^{\delta}_{k})^{*}(y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k}))\rangle$ $\leq 2\lambda_{k}\omega_{k}L\|u_{0}-\hat{u}\|\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|$
((3.11)) $\leq\kappa^{-2}\lambda_{k}^{2}\omega_{k}^{2}L^{2}\|u_{0}-\hat{u}\|^{2}+\kappa^{2}\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2},$
where we have used (3.2) and the inequality $2ab\leq a^{2}+b^{2}$ for $a,b\in\mathbb{R}$ in obtaining 3.2.
By incorporating 3.2, 3.2, (3.3) and (3.2) in 3.2, we attain that
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u^{\delta}_{k}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-\hat{u}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u^{\delta}_{k}-\hat{u}\|\|u_{0}-\hat{u}\|$
((3.12)) $+\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|\big{(}2\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$-\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\big{)}.$
Next, according to the definition of $\hat{k}_{\delta}$ in (3.5), for $k<\hat{k}_{\delta}$ we further derive that
((3.13)) $\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|\leq\frac{1}{\tau}\left(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\frac{M}{k+a}\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\right).$
Using the fact that $\omega_{k}\in[\omega,\Omega],u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ and $\hat{u}\in B_{\rho}(u_{0})$ along with inequality (3.13) in 3.2, we have
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq(1-\lambda_{k})^{2}c(\rho)^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
((3.14)) $-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{2M\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|}{\tau(k+a)}.$
We note that
$(1+\varrho)\big{(}(1-\lambda_{k})^{2}c^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c\big{)}\leq c^{2},$
provided
$c\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}\geq\rho\lambda_{k}(1-\lambda_{k})+\sqrt{(1-\lambda_{k})^{2}+\bigg{(}\frac{1}{1+\varrho}-(1-\lambda_{k})^{2}\bigg{)}(1+\kappa^{-2}\Omega^{2}L^{2})},$
which is true for $c(\rho)$. From (3.6) along with the inequality $2ab\leq a^{2}+b^{2},\forall\hskip 5.69054pta,b\in\mathbb{R}$ and 3.2, we get
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-2\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau}\left(\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}+\frac{1}{(k+a)^{2}}\right).$
$\leq-\left((2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})-(2+M)\omega_{k}\tau^{-1}(1-\lambda_{k})(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
As $\omega_{k}\in[\omega,\Omega],0\leq\lambda_{k}\leq\lambda_{\text{max}}<1$ and $1-\lambda_{k}\leq 1$, we further attain
$\|u^{\delta}_{k+1}-\hat{u}\|^{2}$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y_{i_{k}}^{\delta}-F_{i_{k}}(u^{\delta}_{k})\|^{2}$
$+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
Moreover, given the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we can write the last inequality as
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\|y^{\delta}-F(u^{\delta}_{k})\|^{2}$
$\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
Finally, by taking into consideration the entire conditional, we arrive at
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]$ $\leq-\left((2\omega(1-\lambda_{\text{max}})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})-(2+M)\Omega\tau^{-1}(1+\eta)\right)\mathbb{E}[\|y^{\delta}-F(u^{\delta}_{k})\|^{2}]$
$\hskip 28.45274pt+\frac{c(\rho)^{2}}{1+\varrho}+\frac{M\Omega(1+\eta)}{\tau(k+a)^{2}}.$
This with (3.7), (3.5) for $M=\frac{\varrho c(\rho)^{2}\tau}{\Omega(1+\varrho)(1+\eta)}$ and the inequality $k+a\geq 1$ implies that
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]\leq\frac{c(\rho)^{2}}{1+\varrho}+\frac{\varrho c(\rho)^{2}}{1+\varrho}=c(\rho)^{2}.$
By incorporating Jensen’s inequality (2.2) with $f(x)=x^{2}$, we conclude that
$\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}\leq c(\rho).$
This means that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. Further, by engaging the triangular inequality (2.1) and the fact that $\|u_{0}-\hat{u}\|\leq\rho$, we conclude that
$\mathbb{E}[\|u^{\delta}_{k+1}-u_{0}\|]$ $\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|]+\mathbb{E}[\|\hat{u}-u_{0}\|]\leq\mathbb{E}[\|u^{\delta}_{k+1}-\hat{u}\|^{2}]^{1/2}+\|\hat{u}-u_{0}\|\leq\rho+c(\rho),$
which implies that $u_{k+1}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$.
Next, we will prove that $\hat{k}_{\delta}$ is finite. To establish this, we can utilize 3.2 with (3.13) and (3.7) along with the fact that $1-\lambda_{k}\leq 1$ for $\lambda\in[0,\frac{1}{2})$ to get
$\|u_{k+1}^{\delta}-\hat{u}\|^{2}+D\|y_{i_{k}}^{\delta}-F_{i_{k}}(u_{k}^{\delta})\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}^{\delta}-\hat{u}\|\rho+\frac{M\omega_{k}(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
((3.15)) $\leq\|u_{k}^{\delta}-\hat{u}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\rho^{2}+2\lambda_{k}\rho c(\rho)+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.\hskip 56.9055pt$
Using the inequality that $\lambda_{k}^{2}\leq\lambda_{k}$ and the definition of the measurability of $u_{k}^{\delta}$ along with the additivity of expected values with respect to $\mathcal{F}_{k}$ in 3.2 to get that
((3.16)) $\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+D\|y^{\delta}-F(u_{k}^{\delta})\|^{2}$ $\leq\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}|\mathcal{F}_{k}]+\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}$
$\hskip 56.9055pt+2\rho c(\rho)\lambda_{k}+\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau(k+a)^{2}}.$
Taking the sum of (3.16) from $k=0$ to $k=\hat{k}_{\delta}-1$ to deduce that
((3.17)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1-\lambda_{k}}{(k+a)^{2}}\right).$
From (3.5) and inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})\hskip 5.69054pt\forall\hskip 5.69054pta,b\in\mathbb{R}$, we get
((3.18)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\left(\|F(u_{k}^{\delta})-y^{\delta}\|+\frac{M}{k+a}\right)^{2}\leq 2\left(\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+\frac{M^{2}}{(k+a)^{2}}\right).$
By taking the sum of (3.18) from $k=0$ to $\hat{k}_{\delta}-1$, we get
((3.19)) $\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}\leq 2\left(\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u_{k}^{\delta})-y^{\delta}\|^{2}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right).$
Using (3.17) in (3.19) to get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}$ $\leq 2\left[D^{-1}\left\{\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\sum_{k=0}^{\hat{k}_{\delta}-1}\lambda_{k}\right.\right.$
$\left.\left.\hskip 85.35826pt+\frac{M\Omega(1+\eta)}{\tau}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right\}+M^{2}\sum_{k=0}^{\hat{k}_{\delta}-1}\frac{1}{(k+a)^{2}}\right].$
By applying (3.8) and the result $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$, we further get
$\tau^{2}\|y^{\delta}-y^{\dagger}\|^{2}\hat{k}_{\delta}<\infty.$
Hence $\hat{k}_{\delta}$ is finite. Finally, from the definition of $\hat{k}_{\delta}$ in (3.5), we get
$\frac{M}{\hat{k}_{\delta}+a}\leq\tau\|y^{\delta}-y^{\dagger}\|\to 0\text{\quad as\quad}\delta\to 0.$
Therefore, we must have $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$. ∎
Remark 3.2.
For the definition of $c(\rho)$ to be meaningful, there must exist a positive real root of the quadratic equation
$\left((1-\lambda_{k})^{2}-\frac{1}{1+\varrho}\right)c^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\rho^{2}+2\lambda_{k}(1-\lambda_{k})\rho c=0,$
which is satisfied if the condition $\left((1-\lambda_{k})^{2}-\frac{1}{1+\varrho}\right)<0$ holds. Therefore, a sufficiently small positive constant $\varrho$ is required in (3.6), due to the restriction that $\lambda_{k}\in[0,\frac{1}{2})$.
Lemma 3.2.
Let Assumption 3.1 hold and let the sequence $\{\lambda_{k}\}$ be chosen such that (3.7) and (3.8) hold. Additionally, let $\{y^{\delta}\}$ represent a family of noisy data that fulfill Assumption 3.2. Furthermore, assume that $k_{*}:=k_{*}(y^{\delta})$ is the integer determined by Rule 1. Then
$\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0\quad\text{as}\quad\delta\to 0.$
Consequently, we have
$\|F(u^{\delta}_{k_{*}(y^{\delta})})-y^{\delta}\|\to 0\quad\text{and}\quad k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0\quad\text{as}\quad\delta\to 0.$
Proof.
By employing the condition (3.8) and the fact that $\sum_{k=0}^{\infty}\frac{1}{(k+a)^{2}}\leq\sum_{k=1}^{\infty}\frac{1}{k^{2}}=\frac{\pi^{2}}{6}$ for $a\geq 1$ in (3.17), we arrive at
((3.20)) $\sum_{k=0}^{\hat{k}_{\delta}-1}\|y^{\delta}-F(u^{\delta}_{k})\|^{2}\leq D^{-1}C_{1},$
where $C_{1}=\left(\rho^{2}+\left(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho)\right)\lambda+\frac{M\Omega\pi^{2}(1+\eta)}{6\tau}\right)$. By minimality of $\Psi(k_{*},y^{\delta})$, it follows from Rule 1 that
$\|F(u_{k}^{\delta})-y^{\delta}\|^{2}=\frac{\Psi(k,y^{\delta})}{k+a}\geq\frac{\Psi(k_{*}(y^{\delta}),y^{\delta})}{k+a}.$
This and (3.20) further yields
((3.21)) $\Psi(k_{*}(y^{\delta}),y^{\delta})\sum^{\hat{k}_{\delta}-1}_{k=0}\frac{1}{k+a}\leq\sum_{k=0}^{\hat{k}_{\delta}-1}\|F(u^{\delta}_{k})-y^{\delta}\|^{2}\leq D^{-1}C_{1}.$
Note that
[…]
Plugging the last estimate in (3.21) yields
$\bigg{(}\log\frac{\hat{k}_{\delta}+a}{a}\bigg{)}\Psi(k_{*}(y^{\delta}),y^{\delta})\leq D^{-1}C_{1}.$
Since $\hat{k}_{\delta}\to\infty$ as $\delta\to 0$, we must have $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Further, we note that
$a\|F(u^{\delta}_{k^{*}(y^{\delta})})-y^{\delta}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
Also, Assumption 3.2 implies that
$(k_{*}(y^{\delta})+a)\varkappa^{2}\|y^{\delta}-y^{\dagger}\|^{2}\leq\Psi(k_{*}(y^{\delta}),y^{\delta}).$
Hence, we conclude that $\|F(u_{k_{*}(y^{\delta})}^{\delta})-y^{\delta}\|\to 0$ and $k_{*}(y^{\delta})\|y^{\delta}-y^{\dagger}\|^{2}\to 0$ as $\delta\to 0,$ since $\Psi(k_{*}(y^{\delta}),y^{\delta})\to 0$ as $\delta\to 0$. Thus, result. ∎
The next result demonstrates that IRSGD iterates $u_{k}^{\delta}$ remains in a ball around $u_{0}$ for any $0\leq k<k_{*}$, where $k_{*}$ is defined via Rule 1 (note that (b) of Lemma 3.1 shows the same for $0\leq k<\hat{k_{\delta}}$).
Lemma 3.3.
Let all the assumptions of Lemma 3.2 hold. Additionally, assume that
((3.22)) $D_{1}:=\kappa^{2}-\Omega(1-\lambda_{k})(1-3\eta)+\omega^{2}L^{2}\geq 0.$
Then, if $u^{\delta}_{k}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k_{*}$, it follows that $u^{\delta}_{k+1}\in B_{c(\rho)}(\hat{u})\subset B_{\rho+c(\rho)}(u_{0})$, provided $\delta$ is sufficiently small.
Proof.
For $k=0$, result is trivial. By assuming $u_{k}^{\delta}\in B_{c(\rho)}(\hat{u})$ for any $0\leq k<k^{*},$ we show that $u_{k+1}^{\delta}\in B_{c(\rho)}(\hat{u})$. From 3.2 and the inequalities $\|u_{k}^{\delta}-\hat{u}\|\leq c(\rho)$ and $\|u_{0}-\hat{u}\|\leq\rho$, it follows that
$\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}$
((3.23)) $\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+2\omega_{k}(1-\lambda_{k})(1+\eta)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)(\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2})$
$+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}$
$\leq\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\lambda_{k}^{2}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}+2\lambda_{k}(1-\lambda_{k})\rho c(\rho)$
$+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
Using the fact that $1-\lambda_{k}\leq 1$ and $\lambda_{k}^{2}\leq\lambda_{k}$ for $\lambda\in[0,\frac{1}{2})$, we get
((3.24)) $\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}$ $-\|u_{k}^{\delta}\ -\hat{u}\|^{2}\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\omega_{k}(1-\lambda_{k})(1+\eta)\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|^{2}$
$+(\omega_{k}(1-\lambda_{k})(1+\eta)-\left(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\right)\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|^{2}.$
Further, the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$ implies that
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}-\|u_{k}^{\delta}\ -\hat{u}\|^{2}|\mathcal{F}_{k}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1-\lambda_{k})(1+\eta)\|y^{\dagger}-y^{\delta}\|^{2}$
$\hskip 28.45274pt+\left(\Omega(1-\lambda_{k})(1-3\eta)-\omega^{2}L^{2}-\kappa^{2})\right)\|F(u_{k}^{\delta})-y^{\delta}\|^{2}.$
By considering the conditional in its whole and utilizing $1-\lambda_{k}\leq 1$, we further reach at
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]-\mathbb{E}[\|u_{k}^{\delta}-\hat{u}\|^{2}]$ $\leq(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda_{k}+\Omega(1+\eta)\mathbb{E}[\|y^{\dagger}-y^{\delta}\|^{2}]$
$\hskip 28.45274pt-D_{1}\mathbb{E}[\|F(u_{k}^{\delta})-y^{\delta}\|^{2}].$
The last inequality derives that
((3.25)) $\begin{split}\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]+D_{1}\sum_{l=0}^{k}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]&\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k}\lambda_{l}\\ &\hskip 28.45274pt+\Omega(1+\eta)(k+1)\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\end{split}$
and hence by using (3.8) and the fact that $k<k_{*}$, where $k_{*}$ is a finite integer, we get
$\mathbb{E}[\|u_{k+1}^{\delta}-\hat{u}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
From Lemma 3.2, one can note that for sufficiently small $\delta$, there exist a constant $\xi$ such that
((3.26)) $\Omega(1+\eta)k_{*}(y^{\delta})\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}]\leq\xi,\ \text{with}\ \rho^{2}+\ (\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\lambda+\xi\leq c(\rho)^{2}.$
Hence as a result, we get
$\mathbb{E}[\|u_{k+1}^{\delta}\ -\hat{u}\|^{2}]\leq c(\rho)^{2},$
which completes the proof. ∎
Corollary 3.1.
Let the assumptions of Lemma 3.1 be fulfilled. Moreover assume that IRSGD method (3.1) is stopped according to Rule 1, then for sufficiently small $\delta$, we have
$\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
In particular, for the case $y^{\delta}=y^{\dagger}$ ((((i.e., if $\delta=0)$, we have
$\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty,$
where $D_{1}$ is defined in (3.22).
Proof.
From (3.25), we get
((3.27)) $D_{1}\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{l}+\Omega(1+\eta)k_{*}\mathbb{E}[\|y^{\delta}-y^{\dagger}\|^{2}].$
Using (3.26), we note that
((3.28)) $\sum_{l=0}^{k_{*}-1}\mathbb{E}[\|F(u_{l}^{\delta})-y^{\delta}\|^{2}]\leq D_{1}^{-1}\left((\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{k_{*}-1}\lambda_{k}+\rho^{2}+\xi\right)<\infty.$
Next, for the case $\delta=0$, it follows directly from (3.27) that
((3.29)) $\sum_{l=0}^{\infty}\mathbb{E}[\|F(u_{l})-y^{\dagger}\|^{2}]\leq D_{1}^{-1}\left(\rho^{2}+(\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})+2\rho c(\rho))\sum_{l=0}^{\infty}\lambda_{k}\right)<\infty.$
Here we have used the fact that $\sum_{l=0}^{\infty}\lambda_{k}<\infty.$ Thus, (3.28) and (3.29) complete the proof. ∎
Remark 3.3.
If $\lambda_{\text{max}}=0$, the parameter $\kappa$ in conditions (3.7) and (3.22) can be set to zero. This adjustment mirrors the conditions required to establish the same results for the stochastic gradient descent (SGD) method (1.9).
3.3. Convergence for non noisy case
The following result illustrate that the sequence of IRSGD iterates $\{u_{k}\}_{k\geq 1}$ in (3.1) with exact data $y$ forms a Cauchy sequence.
Lemma 3.4.
Let all the assumptions of Lemma 3.1 hold. Additionally assume that
((3.30)) $D_{2}:=(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2})\geq 0.$
Then the IRSGD method (3.1) for $\delta=0$ produces a sequence of iterates $\{u_{k}\}_{k\geq 1}$ which is a Cauchy sequence.
Proof.
For a solution $u^{\dagger}$ of (1.2), we define $e_{k}:=u_{k}-u^{\dagger}$. As indicated in part (b) of Lemma 3.1, $\mathbb{E}[\|e_{k}\|^{2}]$ is a bounded sequence and hence it has a convergent subsequence $\mathbb{E}[\|e_{k_{n}}\|^{2}]$. Let it converge to some $\epsilon\geq 0$. It follows directly from 3.2 that, for $\delta=0$ we have
$\|u_{k+1}-u^{\dagger}\|^{2}$ $\leq(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\|$
$-(2\omega_{k}(1-\lambda_{k})(1-\eta)-\omega_{k}^{2}L^{2}-\kappa^{2})\|y_{i_{k}}-F_{i_{k}}(u_{k})\|^{2}.$
From the definition of $F$ and the measurability of $u_{k}^{\delta}$ with respect to $\mathcal{F}_{k}$, we deduce that
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}|\mathcal{F}_{k}]$ $\leq\mathbb{E}[(1-\lambda_{k})^{2}\|u_{k}-u^{\dagger}\|^{2}+\lambda_{k}^{2}(1+\kappa^{-2}\omega_{k}^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\|u_{k}-u^{\dagger}\|\|u_{0}-u^{\dagger}\||\mathcal{F}_{k}]$
$-\|y-F(u_{k})\|^{2}(2\omega(1-\lambda_{k})(1-\eta)-\Omega^{2}L^{2}-\kappa^{2}).$
By using (3.30) and considering the full conditional, we reach at
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]+D_{2}\mathbb{E}[\|y-F(u_{k})\|^{2}]$ $\leq(1-\lambda_{k})^{2}\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]+\lambda_{k}^{2}(1+\kappa^{-2}\Omega^{2}L^{2})\|u_{0}-u^{\dagger}\|^{2}$
$+2\lambda_{k}(1-\lambda_{k})\mathbb{E}[\|u_{k}-u^{\dagger}\|]\|u_{0}-u^{\dagger}\|.$
This means that
$\mathbb{E}[\|u_{k+1}-u^{\dagger}\|^{2}]^{1/2}\leq(1-\lambda_{k})\mathbb{E}[\|u_{k}-u^{\dagger}\|^{2}]^{1/2}+\rho\lambda_{k}\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}.$
Hence by induction, we further get that
((3.31)) $\mathbb{E}[\|e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{m}\|^{2}]^{1/2}\prod_{l=m}^{k-1}(1-\lambda_{l})+\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}(1-\prod_{l=m}^{k-1}(1-\lambda_{l}))$
for $m<k,$ where we have used the formula
$1-\prod_{l=m}^{n}(1-\lambda_{l})=\sum_{j=m}^{n}\lambda_{j}\prod_{l=j+1}^{n}(1-\lambda_{l})$
which, under the standard condition that $\prod_{l=n+1}^{n}(1-\lambda_{l})=1$, is satisfied for $n\geq m$.
If $k_{n-1}<k<k_{n}$, then according to (3.31) it follows that
$\mathbb{E}[\|e_{k}\|^{2}]^{1/2}$ $\leq\mathbb{E}[\|e_{k_{n-1}}\|^{2}]^{1/2}+(\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}-\mathbb{E}[\|e_{k_{n-1}}\|^{2}]^{1/2})\left(1-\prod_{l=k_{n-1}}^{k-1}(1-\lambda_{l})\right),$
$\mathbb{E}[\|e_{k}\|^{2}]^{1/2}$ $\geq\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2}-(\rho\sqrt{1+\kappa^{-2}\Omega^{2}L^{2}}-\mathbb{E}[\|e_{k_{n}}\|^{2}]^{1/2})\left(1-\prod_{l=k}^{k_{n-1}}(1-\lambda_{l})\right).$
These estimates, along with the convergence of $\mathbb{E}[\|e_{k_{n}}\|^{2}]$, immediately imply
$\lim_{k\to\infty}\mathbb{E}[\|e_{k}\|^{2}]^{1/2}=\epsilon^{1/2},$
since (3.8) implies that
((3.32)) $\prod_{l=m}^{k-1}(1-\lambda_{l})\to 1\text{ as }m\to\infty,\;k>m.$
Further, to show the convergence of the iterates $u_{k}$, it is crucial to note that $\mathbb{E}[\langle\cdot,\cdot\rangle]$ serves as an inner product. For any $n\geq k$, we select an index $m$ satisfying $n\geq m\geq k$ such that
((3.33)) $\mathbb{E}[\|y-F(u_{m})\|^{2}]\leq\mathbb{E}[\|y-F(u_{i})\|^{2}]\quad\forall k\leq i\leq n.$
Using inequality $\mathbb{E}[\|e_{n}-e_{k}\|^{2}]^{1/2}\leq\mathbb{E}[\|e_{n}-e_{m}\|^{2}]^{1/2}+\mathbb{E}[\|e_{m}-e_{k}\|^{2}]|^{1/2}$ and the identities
$\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{j},e_{m}\rangle]+\mathbb{E}[\|e_{n}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
$\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ $=2\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]+\mathbb{E}[\|e_{k}\|^{2}]-\mathbb{E}[\|e_{m}\|^{2}],$
it is sufficient to establish that $\mathbb{E}[\|e_{n}-e_{m}\|^{2}]$ and $\mathbb{E}[\|e_{m}-e_{k}\|^{2}]$ tend to zero as $k\rightarrow\infty$. As $k\rightarrow\infty$, the last two terms on the right-hand sides of the preceding two identities tend to $\epsilon-\epsilon=0$. This is due to the convergence of $\mathbb{E}[\|e_{k}\|^{2}]$ to $\epsilon$, as discussed above. Next, we proceed to demonstrate that the term $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]$ also approaches to zero as $k\rightarrow\infty$. In fact, by virtue of the definition of $u_{k}$, we have the following representation
$e_{k}=e_{m}\prod_{l=m}^{k-1}(1-\lambda_{l})+\sum_{j=m}^{k-1}\omega_{j}F_{i_{j}}^{\prime}(u_{j})^{\ast}(y_{i_{j}}-F_{i_{j}}(u_{j}))\prod_{l=j+1}^{k-1}(1-\lambda_{l})+(u_{0}-u^{\dagger})\prod_{l=m}^{k-1}(1-\lambda_{l}).$
for $m<k$. Using this formula, we obtain the estimate
$|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y_{i_{j}}-F_{i_{j}}(u_{j}),F_{i_{j}}^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle],$
$|\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]|$ $\leq\left(1-\prod_{l=m}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$
$+\sum_{j=m}^{n-1}\omega_{j}\left(\prod_{l=j+1}^{n-1}(1-\lambda_{l})\right)|\mathbb{E}[\langle y-F(u_{j}),F^{\prime}(u_{j})(u_{m}-u^{\dagger})\rangle]|.$
Since $|\mathbb{E}[\langle u_{0}-u_{m},e_{m}\rangle]|$ is bounded, the first term on the right-hand side tends to zero due to (3.32). Since $1-\lambda_{l}\leq 1$, the second term on the right-hand side can be estimated by
((3.34)) $\Omega(1+3\eta)\sum_{j=m}^{n-1}\mathbb{E}[\|y-F(u_{j})\|^{2}].$
Due to (3.8) and Corollary 3.1, it follows that (3.34) tends to zero as $m\to\infty$. This means $\mathbb{E}[\langle e_{m}-e_{n},e_{m}\rangle]\to 0$. Analogously, we have $\mathbb{E}[\langle e_{m}-e_{k},e_{m}\rangle]\to 0$. Thus, result holds. ∎
In the following theorem, we establish the convergence of non-noisy iterates.
Theorem 3.1.
Suppose that all the assumptions of Lemma 3.1 and (3.30) hold. Then, the IRSGD method (3.1) corresponding to $\delta=0$ generates a sequence of iterates $\{u_{k}\}_{k\geq 1}$ that converges to an element of $\mathcal{D}_{\rho+c(\rho)}(u^{\dagger})$ in $U$, i.e.,
$\lim_{k\to 0}\mathbb{E}[\|u_{k}-\hat{u}\|^{2}]=0,\quad\hat{u}\in\mathcal{D}_{\rho+c(\rho)}(u^{\dagger}).$
Proof.
The proof is similar to [2, Theorem 1]. ∎
3.4. Convergence for noisy case
The next result demonstrates the pathwise stability of the IRSGD iterates $u_{k}^{\delta}$ as $\delta\to 0$. Specifically, for any fixed integer $k\geq 0$ and any sample path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, it holds that $\|u_{k}^{\delta}-u_{k}\|^{2}\to 0$ as $\delta\to 0$. As a result, the expectation $\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]\to 0$ as $\delta\to 0$, confirming the pathwise stability of the iterates for each realization within the filtration $\mathcal{F}_{k}$.
Lemma 3.5.
Let all the assumptions of Lemma 3.1 be satisfied. For any fixed $k\in\mathbb{N}$ and any path $(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}$, let $u_{k}$ and $u_{k}^{\delta}$ be the IRSGD iterates along the path for exact data $y$ and noisy data $y^{\delta}$, respectively. Then
$\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
Proof.
We use induction to prove the assertion. For $k=0$ it is trivial. Suppose that it holds for all indices up to $k$ and any path in $\mathcal{F}_{k}$. From (3.1), for any fixed path $(i_{1},\ldots,i_{k})$, we have
$u_{k+1}^{\delta}-u_{k+1}$ $=(1-\lambda_{k})(u_{k}^{\delta}-u_{k})-\omega_{k}((F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*})(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})$
$+F_{i_{k}}^{\prime}(u_{k})^{*}((F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}))).$
Thus, by the triangle inequality, we get
$\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq(1-\lambda_{k})\|u_{k}^{\delta}-u_{k}\|+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|$
((3.35)) $+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
Next, we claim that for any specific value of $k$, (i) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}\|$ is bounded, and (ii) $\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}_{k}}\|u_{k}^{\delta}-u_{k}\|$ is uniformly bounded. Indeed, by (3.2) and (3.3), we derive that
$\|u_{k+1}-\hat{u}\|$ $\leq\|u_{k}-\hat{u}\|+\omega_{k}\|F_{i_{k}}^{\prime}(u_{k})^{*}\|\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\lambda_{k}\|u_{0}-u_{k}\|$
$\leq\left(1+\lambda_{k}+\omega_{k}\frac{L^{2}}{1-\eta}\right)\|u_{k}-\hat{u}\|+\lambda_{k}\|u_{0}-\hat{u}\|.$
This, together with an induction argument, establishes the claim (i). Similarly, we note that
$\|F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta}\|$ $\leq\|F_{i_{k}}(u_{k}^{\delta})-F_{i_{k}}(u_{k})\|+\|F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger}\|+\|y_{i_{k}}^{\dagger}-y_{i_{k}}^{\delta}\|$
((3.36)) $\leq\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|,$
and consequently 3.4 and 3.4 derive that
$\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq\omega_{k}\left(\frac{L}{1-\eta}\left(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|\right)+\|y^{\delta}-y^{\dagger}\|\right)\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
$+\omega_{k}L\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|+(1-\lambda_{k})\|u_{k}^{\delta}-u_{k}\|.$
This, along with mathematical induction establishes claim (ii). Let $c=\frac{L}{1-\eta}\sup_{(i_{1},\ldots,i_{k-1})\in\mathcal{F}^{k}}(\|u_{k}^{\delta}-u_{k}\|+\|u_{k}-\hat{u}\|)+\|y^{\delta}-y^{\dagger}\|$. Then, it follows from 3.4, 3.4 and (3.2) that
$\lim_{\delta\to 0^{+}}\|u_{k+1}^{\delta}-u_{k+1}\|$ $\leq(1-\lambda_{k})\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|+c\omega_{k}\lim_{\delta\to 0^{+}}\|F_{i_{k}}^{\prime}(u_{k}^{\delta})^{*}-F_{i_{k}}^{\prime}(u_{k})^{*}\|$
$+\omega_{k}L\lim_{\delta\to 0^{+}}\|(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-(F_{i_{k}}(u_{k})-y_{i_{k}}^{\dagger})\|.$
Finally, the result follows from the continuity of the operators $F_{i}$ and $F_{i}^{\prime}$, i.e.,
$\lim_{\delta\to 0^{+}}\|u_{k}^{\delta}-u_{k}\|=0\implies\lim_{\delta\to 0^{+}}\mathbb{E}[\|u_{k}^{\delta}-u_{k}\|^{2}]=0.$
This completes the proof. ∎
The final result of this section establishes the regularizing nature of IRSGD method.
Theorem 3.2.
Let all the assumptions of Lemma 3.1 hold. Let $\{y^{\delta}\}$ be the sequence of noisy data such that Assumption 3.2 hold and let IRSGD iteration corresponding to noisy data is stopped at $k_{*}(y^{\delta})$ defined via Rule 1. Then there exist a solution $\hat{u}$ of (1.2) such that
$\mathbb{E}[\|u_{k_{*}(y^{\delta})}^{\delta}-\hat{u}\|^{2}]\to 0\quad\text{ as}\quad\delta\to 0.$
Proof.
Let $\{y^{\delta_{m}}\}$ be a subsequene of $\{y^{\delta}\}$, and let $K:=\lim\inf_{m\to\infty}k_{*}(y^{\delta_{m}}).$ By taking a subsequence of $\{y^{\delta_{m}}\}$ if necessary, we may assume that $K=\lim_{m\to\infty}k_{*}(y^{\delta_{m}})$ and, according to Lemma 3.5, we have
((3.37)) $\mathbb{E}[\|u_{k}^{\delta_{m}}-u_{k}\|^{2}]\to 0\quad\text{as}\quad m\to\infty.$
for all $0\leq k\leq K.$ Let $\hat{u}$ be the limit of $u_{k}$, which is a solution of (1.2) in $B_{\rho+c(\rho)}(u_{0}).$ First, we assume that $K<\infty,$ then we have $k_{*}(y^{\delta_{m}})=K$ for large $m.$ Thus, from Lemma 3.2 it follows that
$\|F(u_{K}^{\delta_{m}})-y^{\delta_{m}}\|\rightarrow 0\quad\text{as}\quad m\rightarrow\infty.$
By engaging (3.37) and the continuity of $F$, we have $F(u_{K})=y^{\dagger}.$ In other words, $K$th iteration of IRSGD method with exact data is a solution of $F(u)=y^{\dagger}.$ Hence, the iteration terminates with $\hat{u}=u_{K}$ and $u_{K}^{\delta_{m}}\to\hat{u}$ as $m\to\infty.$
If $K=\infty,$ then Theorem 3.1 and Lemma 3.5 guarantees that, for each $\epsilon>0,$ there exist $\hat{k}\in\mathbb{N}$ and $\overline{m}\in\mathbb{N}$ such that
((3.38)) $\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}]<\frac{\epsilon}{8},\quad\text{and}\quad\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}]<\frac{\epsilon}{8}\quad\text{for all}\quad m\geq\overline{m}.$
At this stage, we apply the full conditional on (3.16) to arrive at
((3.39)) $\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq 2\mathbb{E}[\|u_{\hat{k}}^{\delta_{m}}-u_{\hat{k}}\|^{2}+2\mathbb{E}[\|u_{\hat{k}}-\hat{u}\|^{2}+C_{2}\sum_{l=\hat{k}}^{K-1}\lambda_{l}+C_{3}\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}},$
where $C_{2}=2\rho^{2}(1+\kappa^{-2}\Omega^{2}L^{2})$ and $C_{3}=\frac{M\Omega(1-\lambda_{k})(1+\eta)}{\tau}.$ By choosing $\hat{k}$ such that
$\sum_{l=\hat{k}}^{K-1}\lambda_{l}\leq\frac{\epsilon}{4C_{2}}\quad\text{and}\quad\sum_{l=\hat{k}}^{K-1}\frac{1}{(k+a)^{2}}\leq\frac{\epsilon}{4C_{3}},$
and plugging these estimates along with (3.38) in (3.39), we finally get
$\mathbb{E}[\|u_{K}^{\delta_{m}}-\hat{u}\|^{2}]\leq\epsilon\quad\text{for all}\quad m\geq\overline{m}.$
This completes the proof. ∎
4. Convergence rates
The main results of this section are presented in Theorem 4.3 and Theorem 4.4, which respectively address the rate of convergence under the discrepancy principle (1.7) and the rate of convergence under the heuristic rule 1. As a by-product, convergence rates for SGD (1.9) under these two stopping criteria can also be derived by selecting $\lambda_{k}=0$ in the iterative scheme (3.1). These outcomes signify the second major contribution of the work.
Organisation: This section is organised as follows. At the outset, we introduce additional assumptions necessary for the analysis of convergence rates. Subsections 4.1 and 4.2 are dedicated to Theorems 4.1, 4.2, respectively and these theorems play an important role in deriving the convergence rates. In subsection 4.3, we establish convergence rates under a-posteriori stopping criteria, while subsection 4.4 focuses on deriving convergence rates using heuristic stopping criteria.
We employ the following shorthand notation. Let
((4.1)) $K_{i}=F_{i}^{\prime}(u^{\dagger}),\quad K=\begin{pmatrix}K_{0}\\ \vdots\\ K_{P-1}\end{pmatrix}\ \ \text{and}\quad B=K^{*}K=\sum_{i=0}^{P-1}K_{i}^{*}K_{i}.$
Additionally, we commonly use the abbreviated notation
((4.2)) $\prod_{j}^{k-1}(B_{\lambda})=\prod_{i=j}^{k-1}(I-\lambda_{i}-\omega_{i}B),$
with the convention $\prod_{j}^{k}(B_{\lambda})=I$ for $j>k$. Also, for $h\geq 0,$ and $j\in\mathbb{N},$ we define
((4.3)) $\tilde{h}=h+\frac{1}{2}\quad\text{and}\quad\Theta_{j}^{h}=\|B^{h}\prod_{j+1}^{k-1}(B_{\lambda})\|.$
Note that our primary focus is on the cases where $h=0$ and $h=\frac{1}{2}$.
Following the approach in [14], we assume the following conditions.
Assumption 4.1 (Range invariance condition).
There is a family of bounded operators $\{R_{u}^{i}:u\in B_{\rho+c(\rho)}(u_{0})\}$ for all $i\in\{0,1,\ldots,P-1\}$ such that
$F_{i}^{\prime}(u)=R_{u}^{i}F_{i}^{\prime}(u^{\dagger})\quad\text{and}\quad R_{u}=\text{diag}(R_{u}^{i}):Y^{P}\to Y^{P},$
with
$\|R_{u}-I\|\leq c_{R}\|u-u^{\dagger}\|,\quad u\in B_{\rho+c(\rho)}(u_{0})$
for some positive constant $c_{R}>0$ and $R_{u}=\text{diag}(R_{u}^{i})$ (diagonal matrix).
This assumption has been validated for a range of nonlinear inverse problems, such as parameter identification and nonlinear integral equations (see [8]).
Assumption 4.2 (Source condition [16]).
There exist $\gamma\in(0,1/2)$ and $z\in U$ such that
((4.4)) $u^{\dagger}-u_{0}=(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z.$
This assumption reflects a particular smoothness requirement for the exact solution $u^{\dagger}$. The restriction on $\gamma$ arises due to technical considerations (see, for example, (7.9)).
Assumption 4.3 (Stochastic range invariance condition).
Under the notations defined in Assumption 4.1, there exist $\theta\in(0,1]$ and $c_{R}>0$ such that for the function $Q(u)=F(u)-y^{\delta}$ or $Q(u)=F^{\prime}(u^{\dagger})(u-u^{\dagger}),u_{k}^{\delta}\in B_{\rho+c(\rho)}(u_{0})$ and $w_{s}=su_{k}^{\delta}+(1-s)u^{\dagger},s\in[0,1],$ there holds
$\mathbb{E}\left[\|(I-R_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}},$
$\mathbb{E}\left[\|(I-R^{*}_{w_{s}})Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}$ $\leq c_{R}\mathbb{E}\left[\|u^{\delta}_{k}-u^{\dagger}\|^{2}\right]^{\frac{\theta}{2}}\mathbb{E}\left[\|Q(u^{\delta}_{k})\|^{2}\right]^{\frac{1}{2}}.$
This assumption plays a crucial role in the convergence rates analysis (see [14, Lemma 4.6]) and allows for bounding the terms that involves conditional dependence.
4.1. Recursion on the bias
In this subsection, we derive a recursive formula for the bias of the IRSGD iterate $u_{k}^{\delta}.$ The following lemma gives a useful representation of the mean $\mathbb{E}[e_{k}^{\delta}]$.
Lemma 4.1.
Under the conditions of Assumption 4.1, the error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$ exhibits
((4.5)) $\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}),$
with the vector $v_{k}\in Y^{P}$ is given by
((4.6)) $v_{k}=\underbrace{-\left(F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})\right)}_{\text{Linearization error}}+\underbrace{(I-R_{u^{\delta}_{k}}^{*})(F(u^{\delta}_{k})-y^{\delta})}_{\text{Range invariance}}.$
Proof.
From the definition of IRSGD iterates $u_{k}^{\delta}$ in (3.1) we have
$e_{k+1}^{\delta}$ $=(1-\lambda_{k})e_{k}^{\delta}-\omega_{k}F^{\prime}_{i_{k}}(u_{k}^{\delta})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u^{\dagger}-u_{0}).$
The relation $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$ from Assumption 4.1 and (4.1) leads to
$e^{\delta}_{k+1}=(1-\lambda_{k})e^{\delta}_{k}-\omega_{k}\{K^{\ast}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{\ast}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})-K^{\ast}_{i_{k}}v_{k,i_{k}}\}-\lambda_{k}(u^{\dagger}-u_{0}),$
with the random variable $v_{k,i_{k}}$ defined by
((4.7)) $v_{k,i_{k}}=-\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})\right)+(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}}).$
Consequently, by the measurability of $u^{\delta}_{k}$ (and hence $e^{\delta}_{k}$) with regard to $\mathcal{F}_{k}$, we write
$\mathbb{E}[e^{\delta}_{k+1}\mid\mathcal{F}_{k}]$ $=(I-\lambda_{k}-\omega_{k}B)e^{\delta}_{k}-\omega_{k}K^{\ast}(y^{\dagger}-y^{\delta})+\omega_{k}K^{\ast}v_{k}-\lambda_{k}(u^{\dagger}-u_{0}).$
Then taking the full conditional, using (4.2) and applying the recursion repeatedly yields
$\mathbb{E}[e_{k}^{\delta}]=\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)-\sum_{j=0}^{k-1}\lambda_{j}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}).$
This brings the proof to a conclusion. ∎
Remark 4.1.
The term $v_{k}$ in equation (4.6) accounts for both the linearization error of the nonlinear operator $F$, expressed as $F(u^{\delta}_{k})-F(u^{\dagger})-K(u^{\delta}_{k}-u^{\dagger})$, and the range invariance of the derivative $F^{\prime}(u)$ as specified in (ii) of Assumption 3.1 and Assumption 4.1.
The following conclusion provides a reasonable bound on the mean $\mathbb{E}[v_{j}]$.
Lemma 4.2.
Under Assumption 3.1, for $v_{j}$ defined in (4.6), the following inequality holds
((4.8)) $\|\mathbb{E}[v_{j}]\|\leq\frac{(3-\eta)c_{R}}{2(1-\eta)}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\mathbb{E}[\|B^{\frac{1}{2}}e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\frac{1}{2}}\delta.$
Proof.
See [14, Lemma 4.3] for the proof, where $c_{R}>0$ denotes the same constant as specified in Assumption 4.1. ∎
Finally, in the next theorem, we establish a bound for $e_{k}^{\delta}$ in terms of a weighted norm.
Theorem 4.1.
Under Assumptions 3.1, 4.1 and 4.2, for any $h\geq 0,$ there holds
((4.9)) $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|),$
where $\|\mathbb{E}[v_{j}]\|$ is same as given in Lemma 4.2.
Proof.
As $B^{h}$ is a bounded linear operator for $h\geq 0$, by Lemma 4.1, we have
$B^{h}\mathbb{E}[e_{k}^{\delta}]$ $=B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})e_{0}+\sum_{j=0}^{k-1}\omega_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)$
((4.10)) $\hskip 28.45274pt-\sum_{j=0}^{k-1}\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0}):=I+II+III.$
To establish the validity of the theorem, it is sufficient to bound all three terms $I,II$ and $III$. We will address each of these terms individually. From (4.4), we have
((4.11)) $\|I\|=\|B^{h}\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma}z\|\leq\|\Pi^{k-1}_{j=0}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
Next, we estimate the term $II$ as
((4.12)) $\Omega\left\|\sum_{j=0}^{k-1}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})K^{\ast}\left(-(y^{\dagger}-y^{\delta})+\mathbb{E}[v_{j}]\right)\right\|\leq\Omega\sum_{j=0}^{k-1}\|B^{\tilde{h}}\Pi^{k-1}_{j+1}(B_{\lambda})\|\left(\delta+\|\mathbb{E}[v_{j}]\|\right).$
Likewise, we can also estimate the term $III$ using (4.4) as
$\sum_{j=0}^{k-1}\|\lambda_{j}B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})(u^{\dagger}-u_{0})\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma}z\|\leq\sum_{j=0}^{k-1}\lambda_{j}\|\Pi^{k-1}_{j+1}(B_{\lambda})B^{\gamma+h}\|\|z\|.$
Plugging this along with (4.11) and (4.12) in 4.1 together with the notation of $\Theta_{j}^{h}$ concludes the proof. ∎
Remark 4.2.
The weighted norm $\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|$ is valuable because the upper bound presented in Theorem 4.1 incorporates $\mathbb{E}[\|B^{\frac{1}{2}}e_{k}^{\delta}\|^{2}]$ (due to $\|\mathbb{E}[v_{j}]\|$) with $h=\frac{1}{2}$. In the context of linear inverse problems where $R_{u}=I$, $c_{R}=0$ and $\|\mathbb{E}[v_{j}]\|=0$ (due to Lemma 4.2), the recursion simplifies to
$\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|\leq\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\delta,$
where these terms correspond to the approximation error and the data error, respectively.
4.2. Recursion on variance
In this subsection, we address the computational variance $\mathbb{E}[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}]$, which arises from the random index $i_{k}$. The variance was first constrained in terms of the iteration noises $H_{j,1}$ and $H_{j,2}$, which are specified below in (4.14).
Lemma 4.3.
Under Assumption 4.1, the IRSGD iterate $u^{\delta}_{k}$ as outlined in (3.1), satisfies
$\mathbb{E}\left[\|B^{h}(u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}])\|^{2}\right]$ $\leq\Omega^{2}\left(\sum_{j=0}^{k-1}(\Theta^{\tilde{h}}_{j})^{2}\mathbb{E}[\|H_{j,1}\|^{2}]+2\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]\right.$
((4.13)) $\left.\hskip 142.26378pt+\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\Theta^{\tilde{h}}_{i}\Theta^{\tilde{h}}_{j}\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|]\right),$
where the random variables $H_{j,1}$ and $H_{j,2}$, respectively, defined as
((4.14)) $H_{j,1}$ $=(K(u^{\delta}_{j}-u^{\dagger})-K_{ij}(u^{\delta}_{j}-u^{\dagger})\varphi_{ij})+((y^{\dagger}-y^{\delta})-(y^{\dagger}_{i}-y^{\delta}_{i})\varphi_{i_{j}}),$
$H_{j,2}$ $=-\mathbb{E}[v_{j}]+v_{j,i_{j}}\varphi_{i_{j}}.$
Here $v_{k}$ and $v_{k,i}$ are defined in (4.6) and (4.7),respectively, and $\varphi_{i}=(0,\ldots,0,1,0,\ldots,0)$ represents the canonical $i$th basis vector in $\mathbb{R}^{P}$.
Proof.
From the IRSGD iterates as defined in (3.1) along with the Assumption 4.1 that $F^{\prime}_{i_{k}}(u^{\delta}_{k})^{\ast}=(R^{i_{k}}_{u^{\delta}_{k}}F^{\prime}_{i_{k}}(u^{\dagger}))^{\ast}=K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}$, we have
((4.15)) $u_{k+1}^{\delta}=u_{k}^{\delta}-\omega_{k}K^{\ast}_{i_{k}}({R^{i_{k}}_{u^{\delta}_{k}}})^{*}(F_{i_{k}}(u_{k}^{\delta})-y_{i_{k}}^{\delta})-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Rearranging the terms of (4.15) to write it as
((4.16)) $u_{k+1}^{\delta}$ $=u_{k}^{\delta}-\omega_{k}\left(K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})+K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+K_{i_{k}}^{*}\left(F_{i_{k}}(u^{\delta}_{k})-F_{i_{k}}(u^{\dagger})\right.\right.$
$\left.\left.-K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-(I-({R^{i_{k}}_{u^{\delta}_{k}}})^{*})(F_{i_{k}}(u^{\delta}_{k})-y^{\delta}_{i_{k}})\right)\right)-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Further, from the definition of $v_{k,i}$ as in (4.7), we arrive at
((4.17)) $u^{\delta}_{k+1}=u^{\delta}_{k}-\omega_{k}K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})+\omega_{k}K^{*}_{i_{k}}v_{k,i_{k}}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
By the definition of $v_{k}$ in (4.6) and the measurability of $u^{\delta}_{k}$ with respect to $\mathcal{F}_{k}$, we obtain
$\mathbb{E}[u^{\delta}_{k+1}|\mathcal{F}_{k}]=u^{\delta}_{k}-\omega_{k}B(u^{\delta}_{k}-u^{\dagger})-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}v_{k}-\lambda_{k}(u_{k}^{\delta}-u_{0}).$
Taking the full conditional yields
((4.18)) $\mathbb{E}[u^{\delta}_{k+1}]=\mathbb{E}[u^{\delta}_{k}]-\omega_{k}B\mathbb{E}[u^{\delta}_{k}-u^{\dagger}]-\omega_{k}K^{*}(y^{\dagger}-y^{\delta})+\omega_{k}K^{*}\mathbb{E}[v_{k}]-\lambda_{k}\mathbb{E}[u_{k}^{\delta}-u_{0}].$
Thus, subtracting (4.18) from (4.17) shows that $x_{k}:=u^{\delta}_{k}-\mathbb{E}[u^{\delta}_{k}]$ satisfies
((4.19)) $x_{k+1}=(I-\lambda_{k}-\omega_{k}B)x_{k}+\omega_{k}G_{k},$
with $x_{0}=0$ and the iteration noise $G_{k}$ given by $G_{k}=G_{k,1}+G_{k,2}$, where
$G_{k,1}$ $=(B(u^{\delta}_{k}-u^{\dagger})-K^{*}_{i_{k}}K_{i_{k}}(u^{\delta}_{k}-u^{\dagger}))+(K^{*}(y^{\dagger}-y^{\delta})-K^{*}_{i_{k}}(y^{\dagger}_{i_{k}}-y^{\delta}_{i_{k}})),$
$\ G_{k,2}$ $=-(K^{*}\mathbb{E}[v_{k}]-K^{*}_{i_{k}}v_{k,i_{k}}).$
Repeatedly applying the recursion (4.19) with $x_{0}=0$ leads to
$x_{k}=\sum_{j=0}^{k-1}\omega_{j}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j}.$
With the decomposition of $G_{j}=G_{j,1}+G_{j,2}$, we directly obtain
((4.20)) $\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $=\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,1}\rangle]$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]$
$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\mathbb{E}[\langle B^{h}\Pi^{k-1}_{i+1}(B_{\lambda})G_{i,2},B^{h}\Pi^{k-1}_{j+1}(B_{\lambda})G_{j,2}\rangle]:=B_{I}+B_{II}+B_{III}.$
Next, we proceed to simplify the three terms. Noting that $u^{\delta}_{j}$ is measurable with respect to $\mathcal{F}_{j}$, we have $\mathbb{E}[G_{j,1}|\mathcal{F}_{j}]=0.$ Due to this, we claim that
((4.21)) $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle]=0,\quad i\neq j.$
Indeed, for $i>j$, there holds
$\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,1}\rangle|\mathcal{F}_{i}]=\langle B^{h}\mathbb{E}[G_{i,1}|\mathcal{F}_{i}],B^{h}G_{j,1}\rangle=0,$
and taking the full conditional therefore establishes (4.21). Consequently, $B_{I}$ simplifies to
((4.22)) $B_{I}=\Omega^{2}\sum_{j=1}^{k}\mathbb{E}[\|B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,1}\|^{2}].$
Moreover, for $i>j$, a similar reasoning leads to $\mathbb{E}[\langle B^{h}G_{i,1},B^{h}G_{j,2}\rangle]=0,$ and thus
((4.23)) $B_{II}=2\Omega^{2}\sum_{i=1}^{k}\sum_{j=i}^{k}\mathbb{E}[\langle B^{h}\Pi^{k}_{i+1}(B_{\lambda})G_{i,1},B^{h}\Pi^{k}_{j+1}(B_{\lambda})G_{j,2}\rangle].$
Next, we proceed to further simplify $G_{j,1}$ and $G_{j,2}$. From the definitions of $H_{j,1}$ and $H_{j,2}$, and noting that $(K^{*})^{\dagger}$ represents the pseudoinverse of $K^{*}$, we obtain $(K^{*})^{\dagger}G_{j}=H_{j,1}+H_{j,2}$. Thus, by using triangle inequality along with (4.22) and (4.23) in (4.20), we arrive at
$\mathbb{E}[\|B^{h}x_{k}\|^{2}]$ $\leq\Omega^{2}\sum_{j=0}^{k-1}\mathbb{E}[\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|^{2}\|H_{j,1}\|^{2}]$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,1}\|\|H_{j,2}\|]$
$+\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=0}^{k-1}\|B^{h+1/2}\Pi^{k-1}_{i+1}(B_{\lambda})\|\|B^{h+1/2}\Pi^{k-1}_{j+1}(B_{\lambda})\|\mathbb{E}[\|H_{i,2}\|\|H_{j,2}\|].$
This together with the notation of $\Theta_{j}^{h}$ completes the proof. ∎
The next result provides the bounds on $H_{j,1}$ and $H_{j,2}$ defined in (4.14).
Theorem 4.2.
Let Assumptions 3.1 and Assumption 4.3 be fulfilled. Then for any $h\in[0,\frac{1}{2}]$, there holds
$\mathbb{E}$ $\left[\|B^{h}(\mathbb{E}[x^{\delta}_{k}]-x^{\delta}_{k})\|^{2}\right]\leq\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}$
$+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta)\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}$
$+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
Proof.
We recall from [14, Lemma 4.6], with the modification being the exclusion of the scaling factor $n^{1/2}$, that
((4.24)) $\mathbb{E}[\|H_{j,1}\|^{2}]^{1/2}$ $\leq\left(\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+\delta\right),$
$\mathbb{E}[\|H_{j,2}\|^{2}]^{1/2}$ $\leq\left(c_{R}(2+\theta-\eta)\frac{(1+\theta)}{(1-\eta)}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}+c_{R}\delta\right)\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{\theta/2}.$
The conclusion directly follows by using (4.24) in Lemma 4.3. ∎
4.3. Convergence rates using a-posteriori stopping criteria
We recall that our primary objective is to establish the convergence rates using the heuristic stopping criteria (i.e., Rule 1). However, before doing so, in this subsection, we first derive the convergence rates based on the a-posteriori stopping rule (1.7). We emphasize that the convergence analysis for the method with the stopping rule (1.7) follows a similar approach to the one we developed for the heuristic rule in Section 3 with some modification being required due to the change in the stopping rule.
In this subsection, we begin by deriving bounds on the stopping index $k_{\delta}$ and the corresponding error $e_{k}^{\delta}:=u_{k}^{\delta}-u^{\dagger}$.
Lemma 4.4.
Suppose that Assumptions 3.1, 4.1 and 4.2 hold with $\|z\|$ sufficiently small, and let $k_{\delta}$ be the integer specified by (1.7). Also, suppose $y^{\delta}$ satisfies (1.3) and let the weighted parameter sequence $\{\lambda_{k}\}_{k\geq 0}$ be chosen as $\lambda_{k}=(k+1)^{-n}$ with $n>1$, then
((4.25)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma}\quad\text{and}\quad\mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}\|z\|k^{-\gamma-1/2},$
for $0\leq k<k_{\delta},$ $C_{\gamma}$ is a positive constant depending only on $\gamma,\eta,n,\tau$ and $\theta$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
Proof.
Using Bias-variance decomposition, we know that
((4.26)) $\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]=\|B^{h}\mathbb{E}[e_{k}^{\delta}]\|^{2}+\mathbb{E}[\|B^{h}(e_{k}^{\delta}-\mathbb{E}[e_{k}^{\delta}])\|^{2}].$
Using assertions of Theorem 4.1 and Theorem 4.2 in (4.26) with $p_{0}=\frac{(2+\theta-\eta)c_{R}}{(1+\theta)(1-\eta)}$, we have
$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\left(\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+\|\mathbb{E}[v_{j}]\|)\right)^{2}\hskip 85.35826pt$
((4.27)) $+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
By plugging $\|\mathbb{E}[v_{j}]\|$ from Lemma 4.2 in (4.27) and the fact that $\frac{(3-\eta)c_{R}}{2(1-\eta)}<p_{0}$, we get
$\mathbb{E}[\|B^{h}e_{k}^{\delta}\|^{2}]\leq\bigg{(}\Theta_{-1}^{h+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{h+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}(\delta+p_{0}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\mathbb{E}[\|B^{1/2}e^{\delta}_{j}\|^{2}]^{1/2}\hskip 28.45274pt$
$+c_{R}\mathbb{E}[\|e^{\delta}_{j}\|^{2}]^{1/2}\delta)\bigg{)}^{2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{\tilde{h}})^{2}\left(\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+\delta\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{\tilde{h}}\Theta_{j}^{\tilde{h}}\left(\mathbb{E}\left[\|B^{1/2}e_{i}^{\delta}\|^{2}\right]^{1/2}+\delta\right)$
$\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{\tilde{h}}\left(p_{0}\mathbb{E}\left[\|B^{1/2}e_{j}^{\delta}\|^{2}\right]^{1/2}+c_{R}\delta\right)\mathbb{E}\left[\|e_{j}^{\delta}\|^{2}\right]^{\theta/2}\right)^{2}.$
Let $s_{j}=\mathbb{E}[\|e_{j}^{\delta}\|^{2}]$ and $t_{j}=\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]$. Then, by first setting $h=0$ and then $h=1/2$ in the last inequality gives
((4.28)) $s_{k}$ $\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}\delta s_{j}^{1/2})\right)^{2}++2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1/2}\Theta_{j}^{1/2}$
$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2},$
((4.29)) $t_{k}$ $\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\lambda_{j}\Theta_{j}^{1/2+\gamma}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}(\delta+p_{0}s_{j}^{1/2}t_{j}^{1/2}+c_{R}s_{j}^{1/2}\delta)\right)^{2}+2\Omega^{2}\sum_{i=0}^{k-1}\sum_{j=i}^{k-1}\Theta_{i}^{1}\Theta_{j}^{1}$
$\left(t_{j}^{1/2}+\delta\right)\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}+\Omega^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}\left(t_{j}^{1/2}+\delta\right)^{2}+\Omega^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}\left(p_{0}t_{j}^{1/2}+c_{R}\delta\right)s_{j}^{\theta/2}\right)^{2}.$
Because of (3.3) and (1.7), we have
$\tau\delta<\|y^{\delta}-F(u_{k}^{\delta})\|\leq\delta+\frac{1}{1-\eta}\|Ke_{j}^{\delta}\|.$
From the measuribility of $e_{j}^{\delta}$ with respect to $\mathcal{F}_{j}$ and the inequality $\|Ke_{j}^{\delta}\|=\|B^{1/2}e_{j}^{\delta}\|,$ we get
((4.30)) $\delta<\frac{1}{(\tau-1)(1-\eta)}\mathbb{E}[\|B^{1/2}e_{j}^{\delta}\|^{2}]^{1/2}=c_{\tau\eta}t_{j}^{1/2},\quad\text{where}\quad c_{\tau\eta}=\frac{1}{(\tau-1)(1-\eta)}.$
For the sequence of weighted parameter $\{\lambda_{k}\}_{k\geq 0}$, where $\lambda_{k}=(k+1)^{-n}$ and using (4.30) in (4.28) and (4.29) with the notations $c^{\prime}=p_{0}+c_{R}c_{\tau\eta}$ and $c^{\prime\prime}=1+c_{\tau\eta}$ , we have
((4.31)) $s_{k}\leq\left(\Theta_{-1}^{\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1/2}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 71.13188pt$
$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1/2}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1/2})^{2}t_{j}+\Omega^{2}(c^{\prime})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1/2}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2},$
((4.32)) $t_{k}\leq\left(\Theta_{-1}^{1/2+\gamma}\|z\|+\sum_{j=0}^{k-1}\Theta_{j}^{1/2+\gamma}(j+1)^{-n}\|z\|+\Omega\sum_{j=0}^{k-1}\Theta_{j}^{1}[c_{\tau\eta}+c^{\prime}s_{j}^{1/2}]t_{j}^{1/2}\right)^{2}\hskip 42.67912pt$
$+2\Omega^{2}c^{\prime}c^{\prime\prime}\sum_{i=0}^{k-1}\Theta_{i}^{1}t_{i}^{1/2}\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}+\Omega^{2}(c^{\prime\prime})^{2}\sum_{j=0}^{k-1}(\Theta_{j}^{1})^{2}t_{j}+\Omega^{2}(p_{0}+c_{R}c_{\tau\eta})^{2}\left(\sum_{j=0}^{k-1}\Theta_{j}^{1}t_{j}^{1/2}s_{j}^{\theta/2}\right)^{2}.$
We claim that for all $0\leq j<k_{\delta}$, for a certain constant $C_{\gamma}$, there holds
((4.33)) $s_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma}\quad\text{and}\quad t_{j}^{1/2}\leq C_{\gamma}\|z\|j^{-\gamma-1/2}.$
For $j=0$, (4.33) holds trivially. This is because Assumption 4.2 implies that
((4.34)) $\|e_{0}\|\leq\|(F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger}))^{\gamma}z\|\leq\|F^{\prime}(u^{\dagger})^{*}F^{\prime}(u^{\dagger})\|^{\gamma}\|z\|\leq\|F^{\prime}(u^{\dagger})\|^{2\gamma}\|z\|.$
Using Assumption 3.1 $(ii)$ in (4.34), we conclude that
$\|e_{0}\|\leq L^{2\gamma}\|z\|\leq C_{\gamma}\|z\|.$
for sufficiently large $C_{\gamma}$. For $j>0$, the proof proceeds by induction. We assume that (4.33) is true for all $0\leq j<k$ for some $k<k_{\delta}$, and then verify it for $j=k$. So using the induction hypothesis in (4.31) and (4.32) along with Lemma A1 and A3 with appropriate choices of $a,b$ and $c$, we have
((4.35)) $s_{k}\leq\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}k^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-\gamma-n}\right)\|z\|$
$+\Omega c_{\tau\eta}\hslash\widetilde{C}_{1/2,1/2,0}(k/2)^{-\gamma}+\Omega c^{\prime}\hslash^{2}\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-2\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{2+\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma(2+\theta)}$
$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\hslash^{2}\left(\frac{1}{e\omega}(l_{1,0}+2\ln k)+2^{-2\gamma-1}\|B\|\right)(k/2)^{-2\gamma-1}+\Omega^{2}(c^{\prime})^{2}\hslash^{2(1+\theta)}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma(1+\theta)}.$
Further, using the fact that for any $r>0,$ there holds
((4.36)) $s^{-r}\ln{s}\leq(er)^{-1}\quad\forall s\geq 0,$
and thus $k^{-1}\ln{k}\leq e^{-1},$ we further obtain from (4.35) using the notation $\hslash=C_{\gamma}\|z\|$ that
((4.37)) $s_{k}\leq\hslash^{2}\bigg{\{}\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}2^{-\gamma}+\left(\left(\frac{\gamma}{e\omega}\right)^{\gamma}\left(\frac{1}{1-\gamma}+\frac{1}{1-n}\right)+2^{1-\gamma-n}\|B^{a}\|\right)(k/2)^{1-n}\right)C_{\gamma}^{-1}$
$+\Omega c_{\tau\eta}\widetilde{C}_{1/2,1/2,0}+\Omega c^{\prime}\hslash\widetilde{C}_{1/2,1/2,1/2}(k/2)^{-\gamma}\bigg{\}}^{2}+2\Omega^{2}c^{\prime}c^{\prime\prime}\hslash^{\theta}\widetilde{C}_{1/2,1/2,0}\widetilde{C}_{1/2,1/2,\theta/2}(k/2)^{-\gamma\theta}$
$+\frac{\Omega^{2}}{2}(c^{\prime\prime})^{2}\bigg{\{}\left(\frac{l_{1,0}}{e\omega}+2^{-2\gamma-1}\|B\|\right)(k/2)^{-1}+\frac{4}{e^{2}\omega}\bigg{\}}+\Omega^{2}(c^{\prime})^{2}\hslash^{2\theta}\widetilde{C}_{1/2,1/2,\theta/2}^{2}(k/2)^{-2\gamma\theta}\Bigg{]}(k/2)^{-2\gamma}$
$:=\hslash^{2}B_{\gamma,k}(k/2)^{-2\gamma}.\hskip 256.0748pt$
At this point, let us assume that $B_{\gamma,k}>0$ be such that $B_{\gamma,k}\leq 2^{-2\gamma}$ (for sufficiently small $\|z\|$ and sufficiently large $C_{\gamma}$, this condition is satisfied), which means that
((4.38)) $s_{k}\leq\hslash^{2}2^{-2\gamma}\left(\frac{k}{2}\right)^{-2\gamma}\leq\hslash^{2}k^{-2\gamma}=C_{\gamma}^{2}\|z\|^{2}k^{-2\gamma}.$
This completes the induction step and the proof for $s_{k}$. The rest part of the proof is moved to Appendix as it follows on the lines of that of $s_{k}$. ∎
Theorem 4.3.
Under the assumptions of Lemma 4.4, we have
((4.39)) $k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}}$
where $C_{\gamma,\eta,n,\tau,\theta}$ is a positive constant depending only on $\gamma,\eta,n,\tau,\theta$. In addition to that
((4.40)) $\mathbb{E}[\|e_{k_{\delta}}^{\delta}\|^{2}]^{1/2}\leq C_{\gamma}C_{\gamma,\eta,n,\tau,\theta}\|z\|^{1/(2\gamma+1)}\delta^{2\gamma/(2\gamma+1)},$
[…]
Proof.
We apply (4.30) along with (4.33) with $k=k_{\delta}$ to obtain
$k_{\delta}^{\gamma+1/2}\leq c_{\tau\eta}C_{\gamma}\|z\|/\delta$
with $C_{\gamma}$ as in Lemma 4.4. Hence, we obtain
$k_{\delta}\leq C_{\gamma,\eta,n,\tau,\theta}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta}=(c_{\tau\eta}C\gamma)^{\frac{2}{2\gamma+1}}.$ The error estimate (4.40) originates from (4.25). ∎
4.4. Convergence rates using heuristic type stopping criteria
In this section, we will derive an error estimate for the expression $u^{\delta}_{k_{*}(y^{\delta})}-u^{\dagger}$ in mean squared norm.
Lemma 4.5.
Let Assumption 3.1 hold and $y^{\delta}$ be a noisy data satisfying Assumption 3.2, also let $k_{*}:=k_{*}(y^{\delta})$ be the integer defined by Rule 1. Additionally, assume that
((4.41)) $\|F(u_{0})-y^{\delta}\|>\tau\|y^{\dagger}-y^{\delta}\|.$
If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.42)) $(k_{*}+a)\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\|z\|^{\frac{2}{2\gamma+1}}\delta^{-\frac{2}{2\gamma+1}},$
where $C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}=\frac{\tau^{2}(1+a)C_{\gamma,\eta,n,\tau,\theta}}{\varkappa^{2}}$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
By Assumption 3.2 and the definition of $k_{*}$ and $k_{\delta}$, we have
$\varkappa^{2}(k_{*}+a)\delta^{2}\leq\Psi(k_{*},y^{\delta})\leq\Psi(k_{\delta},y^{\delta})=(k_{\delta}+a)\|F(u_{k_{\delta}}^{\delta})-y^{\delta}\|^{2}\leq\tau^{2}(k_{\delta}+a)\delta^{2}.$
This implies that
$\varkappa^{2}(k_{*}+a)\leq\tau^{2}(k_{\delta}+a)\leq\tau^{2}(1+a)k_{\delta},\ \text{as}\ k_{\delta}\geq 1$
since $\|F(u_{0})-y^{\delta}\|>\tau\delta$. Therefore, by virtue of Theorem 4.3, the proof is complete. ∎
Lemma 4.6.
Let Assumption 3.1 hold and let $k_{*}$ be the integer defined by Rule 1 and let $y^{\delta}$ satisfy (1.3). If Assumption 4.1 holds with $20<\gamma<1/$ 2 for sufficiently small $\|z\|,$ then
((4.43)) $\mathbb{E}[\|e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma},\ \ \ \text{and}\ \ \ \mathbb{E}[\|B^{1/2}e_{k}^{\delta}\|^{2}]^{1/2}\leq\tilde{C}_{\gamma}\|z\|k^{-\gamma-1/2}$
for $0\leq k\leq k_{*},$ $\tilde{C}_{\gamma}>0$ is a constant depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$. In case of exact data $(\delta=0)$, (4.25) holds for all $k\geq 0.$
Proof.
The proof follows on the similar lines as given in Lemma 4.4 with a small modification in (4.30) due to heuristic type stopping rule $k_{*}$. By virtue of Lemma 4.5, we have
((4.44)) $\delta\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|(k_{*}+a)^{-\gamma-1/2}\leq C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma+1/2}\|z\|j^{-\gamma-1/2}.$
Using (4.44) in (4.28) and (4.29) and following the same procedure as in Lemma 4.4 we get the desired result. ∎
Remark 4.3.
In comparison to Lemma 4.4, the only difference in Lemma 4.6 is a constant factor. This variation arises due to (4.44), which is applied in the proof in place of (4.30).
Theorem 4.4.
Under the assumptions of Lemma 4.6 we have
((4.45)) $\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}},$
with some constants $\tilde{C}_{\gamma},C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}>0,$ depending only on $\gamma,\eta,n,\tau,\theta,\varkappa$ and $a$.
Proof.
We apply (4.42) along with (4.43) with $k=k_{*}$ as
$\mathbb{E}[\|e_{k_{*}}^{\delta}\|^{2}]^{1/2}$ $\leq\tilde{C}_{\gamma}\|z\|C^{-\gamma}_{\gamma,\eta,n,\tau,\theta,\varkappa,a}\left(\frac{\|z\|}{\delta}\right)^{\frac{-2\gamma}{2\gamma+1}}\leq\frac{\tilde{C}_{\gamma}}{C_{\gamma,\eta,n,\tau,\theta,\varkappa,a}^{\gamma}}\|z\|^{\frac{1}{2\gamma+1}}\delta^{\frac{2\gamma}{2\gamma+1}}.$
Hence we get the desired result.
4.5. Comparison of convergence rates
In this subsection, we present a detailed comparison of various iterative methods. As shown in Table 1, the convergence rates achieved by our approach are comparable to those reported in the existing literature. Furthermore, we selected the step size $\omega_{k}\in[\omega,\Omega]$ in such a way that it can either remain constant or decay monotonically, depending on the specific method used.
Method Step-size Stopping criteria Convergence Rate
LM [8] Constant a posteriori (1.7) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma>0$
IRLM [16] Constant a-priori $O(\delta^{1/2})$
SGD [14] Decaying $k^{*}=\left[(\frac{\delta}{\|w\|})^{-\frac{2}{(2\gamma+1)(1-\alpha)}}\right]$ $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$ and $\alpha$ close to $1$
SGD [12] Decaying a posteriori (1.7) Not available
IRSGD (This paper) Any a posteriori (1.7) $O(\delta^{\frac{2\gamma}{(2\gamma+1)}}),\gamma\in(0,\frac{1}{2})$
IRSGD (This paper) Any heuristic type (Rule 1) $O(\delta^{\frac{2\gamma}{2\gamma+1}}),\gamma\in(0,\frac{1}{2})$
Table 1. Convergence rates of various iterative methods.
5. Numerical Experiments
The objective of this section is to discuss the applicability of proposed scheme using Rule 1 on linear as well as nonlinear ill-posed inverse problems.
5.1. Linear Case
Consider the following problem of linear ill-posed system (see [15] for details on this problem)
[…]
derived from first-kind linear integral equations on $[a,b]$ by sampling at $t_{i}\in[a,b]$ for $i=0,\ldots,P-1,$ where $t_{i}=a+\frac{i(b-a)}{P-1}$ for $i=0,\ldots,P-1,$ and the kernel $k(\cdot,\cdot)$ is continuous on $[a,b]\times[a,b]$. In this case, the iterative scheme (3.1) becomes
[…]
where $i_{k}$ is the randomly chosen index at the $k$th iteration step from the index set $\{0,\ldots,P-1\}$. We consider the system (5.1) with $[a,b]=[-6,6],P=1000,$ and $k(t,s)=\varphi(t-s),$ where $\varphi(t)=(1+\cos(\frac{\pi t}{3}))_{\chi_{\{|t|<3\}}}$ to test the effectiveness of technique (5.2). We consider the following solution to be reconstructed
$u^{\dagger}(s)=\sin\left(\frac{\pi s}{12}\right)+\sin\left(\frac{\pi s}{3}\right)+\frac{1}{200}s^{2}(1-s).$
Instead of the exact data $y_{i}=F_{i}(u^{\dagger})$, we use the noisy data
((5.3)) $y^{\delta}_{i}=y_{i}+\delta_{\text{rel}}\times|y_{i}|\times{\epsilon_{i}},\quad i=0,\ldots,P-1,$
where $\epsilon_{i}$ is a random noise that follows a standard Gaussian distribution $N(0,1)$, and $\delta_{\text{rel}}$ is the relative noise level. Based on the division of $[-6,6]$ into $P-1$ subintervals of similar length, the trapezoidal rule approximates the integrals involved in the procedure and discretized into a linear system of size $n=1000$ with the forward operator $F_{n\times n}$. We execute the method (5.2) with the numerical settings as follows:
  • (a)
    Initial guess $u_{0}=0.$
  • (b)
    The step size is chosen as $\omega_{k}=\omega=10^{-3}.$
  • (c)
    The sequence of weighted parameters $\{\lambda_{k}\}_{k\geq 1}$ is chosen with $\lambda_{k}=\frac{1}{k^{2}}.$
  • (d)
    The maximum number of iterations is set to $k_{\text{max}}=1000$.
  • (e)
    We employ noisy data, where $\delta_{\text{rel}}=10^{-1},10^{-2},10^{-3}$ and $10^{-4}$ represent four different relative noise levels.
We mainly perform the three experiments as:
5.1.1. Convergence behavior of $\Psi(k,y^{\delta})$
In order to investigate the convergence behavior of $\Psi(k,y^{\delta})$ as $\delta\to 0$, we present the plots of the values of $\Psi(k,y^{\delta})$ versus $k$ in Figure 1 using the fixed constant $a=100$ (as in Rule 1) and for different levels of noise.
Figure 1. Convergence behavior of $\Psi(k,y^{\delta})$ as defined in Rule 1, alongside the increase in the stopping index $k_{*}$ as the noise level decreases.
Figure 1 clearly shows that the noise level in the input data has a significant effect on the convergence behavior of $\Psi(k,y^{\delta})$. As $\delta_{\text{rel}}$ decreases from $10^{-1}$ to $10^{-4}$, convergence becomes both faster and more stable. Additionally, it demonstrates that $\Psi(k_{*},y^{\delta})\to 0$ as $\delta\to 0$, which supports the assertion in Lemma 3.2. Furthermore, it is also evident that as the noise level decreases, the stopping index $k_{*}$ increases.
5.1.2. Reconstructed solutions
We reconstructed the solution using (5.2) at various noise levels, as shown in Figures LABEL:M=0.1-LABEL:M=0.001. These results illustrate the impact of the constant $a$ (as defined in Rule 1) on reconstruction accuracy. At a noise level of $10^{-1}$, smaller values of $a$ lead to early stopping of the algorithm, resulting in an inaccurate reconstruction. As $a$ increases (e.g., $a=100$ and $a=1000$), the reconstruction accuracy improves significantly.
From Figures LABEL:M=0.01 and LABEL:M=0.001 observe that, as the noise level decreases, the inaccuracy decreases for small values of $a$ (e.g., $a=1$ and $a=10$). However, for $a=1000$, the number of iterations increases significantly compared to the other cases, leading to a higher computational burden. Therefore, for lower noise levels, smaller values of $a$ should be chosen to reduce computational cost and achieve faster results.
Tables 2 provide detailed numerical results (stopping index $k_{*}$ and the relative errors $E^{\delta}_{k}:=\frac{\|u^{\delta}_{k}-u^{\dagger}\|_{L^{2}(\Omega)}}{\|u^{\dagger}\|_{L^{2}(\Omega)}}$) for Rule 1 at various noise levels.
$\delta_{\text{rel}}$ $a$ in Rule 1
[…]
$k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $k_{*}$ $E_{k_{*}}^{\delta}$
$1\text{e}^{-1}$ 18 $3.550\text{e}^{-2}$ 32 $1.565\text{e}^{-2}$ 49 $1.441\text{e}^{-2}$ 249 $4.213\text{e}^{-3}$
$1\text{e}^{-2}$ 133 $4.817\text{e}^{-3}$ 80 $8.626\text{e}^{-3}$ 506 $1.300\text{e}^{-3}$ 565 $1.664\text{e}^{-3}$
$1\text{e}^{-3}$ 172 $2.293\text{e}^{-3}$ 260 $4.982\text{e}^{-3}$ 2044 $5.845\text{e}^{-4}$ 3404 $4.982\text{e}^{-4}$
$1\text{e}^{-4}$ 3304 $5.341\text{e}^{-4}$ 3167 $5.134\text{e}^{-4}$ 867 $6.111\text{e}^{-4}$ 4874 $3.185\text{e}^{-4}$
Table 2. Results of IRSGD with different values of $a$ in Rule 1.
5.1.3. Comparison between SGD and IRSGD
We compare the results of IRSGD with SGD. For this comparison, we set the constant $a$ in the rule (referred to as Rule 1) to $a=100$. The numerical results for the comparison are presented in Table 3.
Methods $\delta_{\text{rel}}$ $k_{*}$ $E_{k_{*}}^{\delta}$ $\Psi(k_{*},y^{\delta})$
IRSGD $1e^{-1}$ 68 $1.58e^{-2}$ 139237.06
SGD $1e^{-1}$ 66 $3.34e^{-2}$ 153135.32
IRSGD $1e^{-2}$ 180 $2.49e^{-3}$ 24514.94
SGD $1e^{-2}$ 192 $1.44e^{-3}$ 27692.34
IRSGD $1e^{-3}$ 258 $5.53e^{-3}$ 20785.89
SGD $1e^{-3}$ 772 $3.40e^{-4}$ 8772.18
IRSGD $1e^{-4}$ 3492 $7.93e^{-4}$ 21809.47
SGD $1e^{-4}$ 4122 $4.59e^{-4}$ 12432.96
IRSGD $1e^{-5}$ 3468 $6.24e^{-4}$ 17168.44
SGD $1e^{-5}$ 3529 $7.71e^{-4}$ 17952.29
Table 3. Comparison of IRSGD and SGD with Rule 1.
We observe that, as noise level decreases (moving down the table), both methods generally require more iterations ($k_{*}$) to achieve an acceptable solution. Lower noise levels yield lower relative errors $E_{k_{*}}^{\delta}$, which is expected as there is less perturbation in the data, allowing both methods to approximate the true solution more accurately. For different noise levels, IRSGD tends to converge in fewer iterations than SGD, indicating that the additional damping term in IRSGD can speed up convergence. However, for very low noise ($\delta_{\text{rel}}=1\times 10^{-5}$), the iteration counts for IRSGD and SGD are similar, suggesting that both methods stabilize in a similar time frame.
5.2. Nonlinear case
This subsection presents the mathematical model that illustrates the Schlieren tomography and validates the proposed methods. The Schlieren optical system produces a light intensity pattern that is proportional to the square of the line integral of pressure along the light path as it passes through the tank (see [10] for more details).
For each $i$ between 00 and $P-1$, let $\sigma_{i}\in S^{1}$ be a set of recording directions such that $\zeta_{i}=\zeta(\theta_{i})=(\cos\theta_{i},\sin\theta_{i})$. The problem is to reconstruct a function $u$ supported on a bounded domain $D:=[-1,1]\times[-1,1]\subset\mathbb{R}^{2}$ from the system of operator equations
$F_{i}(u)=y_{i}^{\delta},\quad\text{for }i=0,\ldots,P-1,$
where the Schlieren transform $F_{i}=R_{i}^{2}$, with the operator $R_{i}$ denoting the Radon transform along the direction $\zeta_{i}$, satisfying $R_{i}:C_{0}^{\infty}(D)\to C_{0}^{\infty}(I)$:
[…]
From [10], we know that each operator $F_{i}:H_{0}^{1}(D)\to L^{2}(I)$ with $I=[-1,1]$ is continuous and Fréchet differentiable with
$F_{i}^{\prime}(u)h=2R_{i}(u)R_{i}(h),\quad\forall h\in H_{0}^{1}(D).$
In addition, the adjoint $F_{i}^{\prime}(u)^{*}:L^{2}(I)\to H_{1}^{0}(D)$ is given by
$F_{i}^{\prime}(u)^{*}g=(I-\Delta)^{-1}\left(2R_{i}^{*}(R_{i}(u)g)\right),\quad\forall g\in L^{2}(I),$
where $R_{i}^{*}:L^{2}(I)\to L^{2}(D)$ is the adjoint of $R_{i}$ given by $(R_{i}^{*}g)(u)=g(u\cdot\sigma_{i}),$ $I$ refers to the identity operator and $\Delta$ represents the Laplace operator acting on $H_{0}^{1}(D)$. The computation of the adjoint operator requires solving an elliptic boundary value problem, which is approximated using a multigrid method implemented through finite difference discretization.
In this numerical experiment, we set the parameters as follows: $\delta_{\text{rel}}=10^{-2}$, $k_{\text{max}}=750$, $\omega_{k}=\omega=5\times 10^{-3}$, and $\lambda_{k}=\frac{1}{k^{3}}$. The true image and initial guess are shown in Figure LABEL:initial_and_true_figures. Using the proposed method, we reconstructed the image with a stopping criterion based on Rule 1. We evaluated the reconstruction quality by tracking both the relative errors and the progression of $\Psi(k,y^{\delta})$ for different values of $a$: specifically, $a=10^{3}$ and $10^{4}$. The results for each case are displayed in Figures LABEL:RE=1000-LABEL:RE=_10000, respectively.
6. Conclusion
In this paper, we present an Iteratively Regularized Stochastic Gradient Descent (IRSGD) method specifically developed to address systems of ill-posed inverse problems, utilizing the Hanke-Raus heuristic rule. Notably, the proposed IRSGD method is fully data-driven, as the Hanke-Raus rule operates independently of any prior knowledge of the noise level. Our convergence analysis is based on the tangential cone condition and assumes Lipschitz continuity of the operators involved.
Furthermore, we have conducted a thorough analysis of the convergence of our method in the mean squared norm. Additionally, we have established convergence rates under both a-posteriori and heuristic rules by leveraging the source condition, range invariance condition, and stochastic range invariance condition for arbitrary step sizes. To demonstrate the practical effectiveness of our approach, we have provided extensive numerical experiments on both linear and nonlinear inverse problems.
Lastly, we outline several directions for future research. First, it is essential to validate our assumptions and theoretical analysis in the context of specific nonlinear inverse problems to further demonstrate the practical applicability of the IRSGD method. Second, following the work in [2, 4, 5, 19], the IRSGD approach can be extended to handle nonsmooth nonlinear problems, thereby broadening its range of applications. Finally, it is important to investigate adaptive step size rules to provide practical insights and guidelines for real-world implementation. The heuristic rule (based on the modified discrepancy principle), which is used in this analysis can be further modified to study the stochastic analysis of regularization techniques presented in [5, 16, 18, 19, 20].
Acknowledgments
Harshit Bajpai would like to express gratitude to the Science and Engineering Research Board (SERB), India, for providing his PhD fellowship through grant CRG/2022/005491. Ankik Kumar Giri also extends his thanks to SERB for supporting his research under the project “Fast and efficient iterative methods for solving nonlinear inverse problems” funded by grant CRG/2022/005491.